### v3

# VMarker-Pro: Probabilistic 3D Human Mesh Estimation from Virtual Markers

Abstract
Monocular 3D human mesh estimation faces challenges due to depth ambiguity and the complexity of mapping images to complex parameter spaces. Recent methods propose to use 3D poses as a proxy representation, which often lose crucial body shape information, leading to mediocre performance. Conversely, advanced motion capture systems, though accurate, are impractical for markerless wild images. Addressing these limitations, we introduce an innovative intermediate representation as virtual markers, which are learned from large-scale mocap data, mimicking the effects of physical markers. Building upon virtual markers, we propose VMarker, which detects virtual markers from wild images, and the intact mesh with realistic shapes can be obtained by simply interpolation from these markers. To address occlusions that obscure 3D virtual marker estimation, we further enhance our method with VMarker-Pro, a probabilistic framework that generates multiple plausible meshes aligned with images. This framework models the 3D virtual marker estimation as a conditional denoising process, enabling robust 3D mesh estimation. Our approaches surpass existing methods on three benchmark datasets, particularly demonstrating significant improvements on the SURREAL dataset, which features diverse body shapes. Additionally, VMarker-Pro excels in accurately modeling data distributions, significantly enhancing performance in occluded scenarios. Code and models are available at https://github.com/ShirleyMaxx/VMarker-Pro.
Index Terms:
Computer vision, 3D human mesh estimation, probabilistic modeling
Figure 1: Top: Mesh estimation results on examples with different body shapes. Pose2Mesh [1] which uses 3D skeletons as the intermediate representation fails to predict accurate shapes. Our virtual marker-based method VMarker obtains accurate estimates. Bottom: VMarker may fail when occlusion happens, e.g.  the occluded left arm is wrongly estimated. To handle the ambiguity, we propose a probabilistic framework VMarker-Pro which could generate multiple reasonable estimates aligning with the image.
1 Introduction
Monocular 3D human mesh estimation aims to estimate 3D positions of the mesh vertices that are on the body surface from an RGB image. The task has attracted a lot of attention from the computer vision and computer graphics communities [1, 2, 3, 4, 5, 6, 7, 8] because it can benefit many applications such as virtual reality [9]. Recently, deep learning-based methods [6, 1, 10] have significantly advanced the accuracy on benchmark datasets.
Pioneer methods [11, 6] propose to regress the pose and shape parameters of the mesh models such as SMPL [12] directly from images, learning a one-to-one mapping from images to the parameter space. While straightforward, their performance usually suffers from several limitations. Firstly, the mapping from the image to the $SO(3)$ parameter space is highly non-linear and suffers from image-model misalignment [10]. Secondly, existing mesh datasets [13, 14, 15] are small and limited to simple laboratory environments due to the complex capturing process. The lack of sufficient training data severely limits its performance. Lastly, due to the inherent depth ambiguity in monocular estimation, multiple 3D meshes may correspond to the same 2D projection. As a result, these deterministic methods may produce incorrect estimates in the presence of occlusions.
Recently, some works [16, 17] begin to formulate mesh estimation as a dense 3D keypoint detection task inspired by the success of volumetric pose estimation [18, 19, 20, 21, 22, 23]. For example, in [16, 17], the authors propose to regress the 3D positions of all vertices. However, it is computationally expensive because it has more than several thousand vertices. Moon et al.  [17] improves the efficiency by decomposing the 3D heatmaps into multiple 1D heatmaps at the cost of mediocre accuracy. Choi et al.  [1] propose to first detect a sparser set of skeleton joints in the images, from which dense 3D meshes are regressed by exploiting the mesh topology. The methods along this direction have attracted increasing attention [1, 10, 24] due to two reasons. First, the proxy task of 3D skeleton estimation can leverage the abundant 2D pose datasets which notably improves the accuracy. Second, mesh regression from the skeletons is efficient. However, important information about the body shapes is lost in extracting the 3D skeletons, which is largely overlooked. As a result, different types of body shapes, whether lean or obese, cannot be accurately estimated, e.g.  the results of Pose2Mesh [1] in Figure 1.
The professional marker-based motion capture method MoSh [5] places physical markers on the body surface to extract meshes with accurate shapes under non-rigid motions. However, the physical markers limit the approach in laboratory environments. We are inspired to think whether we can identify a set of landmarks on the mesh as virtual markers, e.g.  , elbow and wrist, that can be detected from wild images, allowing us to recover accurate body shapes. The desired virtual markers should satisfy several requirements. First, the number of markers should be much smaller than that of the mesh vertices so that we can use volumetric representations to estimate their 3D positions efficiently. Second, the markers should capture the mesh topology so that the intact mesh can be accurately regressed from them. Third, the virtual markers have distinguishable visual patterns and can be detected from images.
In this work, we present a learning algorithm based on archetypal analysis [25] to identify a subset of mesh vertices as the virtual markers that try to satisfy the above requirements to the best extent. Figure 2 shows that the learned virtual markers coarsely outline the body shape and pose which paves the way for estimating meshes with accurate shapes. Then we present a simple framework VMarker for 3D mesh estimation on top of the representation as shown in Figure 3 (top). It first detects the 3D positions of the virtual markers [18]. Then we recover the intact mesh simply by interpolating them. The interpolation weights are pre-trained in the representation learning step and will be adjusted by a light network based on the prediction confidences of the virtual markers for each image.
Although the virtual marker representation eases the mesh estimation task, the vulnerability of deterministic methods to depth ambiguity and occlusions underscores the need for probabilistic modeling. For example, in Figure 1 (bottom), VMarker wrongly estimates the occluded left arm. Expanding upon VMarker, we further develop VMarker-Pro, a probabilistic framework capable of estimating multiple plausible 3D meshes that are consistent with the given 2D observation, as depicted in Figure 3 (bottom). Motivated by recent breakthroughs in generative models, we utilize diffusion models [26] to enhance the diversity and robustness of our mesh predictions. Concretely, we formulate the estimation of 3D virtual markers as a conditional reverse denoising process. Based on the input 2D image and the predicted virtual markers from VMarker, we probabilistically model the 3D marker positions and finally obtain multiple reasonable 3D mesh estimates. For instance, VMarker-Pro could estimate a more naturally swinging left arm during walking than VMarker as shown in the bottom of Figure 1.
We extensively evaluate our approaches on three benchmark datasets. Both VMarker and VMarker-Pro consistently outperforms the state-of-the-art methods across the benchmarks. Notably, VMarker demonstrates remarkable gains on the SURREAL dataset [27], which has a variety of body shapes. Ablation studies further affirm the merits of virtual marker representation, particularly in estimating accurate shapes. Furthermore, VMarker-Pro stands out in its precise modeling of data distributions, markedly enhancing performance in occlusion-heavy scenarios. Finally, experiments show decent generalization ability of our methods which generate visually appealing results for wild images.
A preliminary version of this work was presented in CVPR 2023 [28]. Our new contributions are three-fold. First, we propose VMarker-Pro, which incorporates probabilistic modeling to handle the inherent ambiguity in monocular 3D mesh estimation. This yields a more flexible and robust human mesh distribution estimation. Second, VMarker-Pro outperforms state-of-the-arts (SOTA) on multiple benchmark datasets in terms of the accuracy of distribution. Finally, VMarker-Pro significantly improves the performance in scenarios with occlusions than VMarker and other SOTA methods, demonstrating its ability to resolve ambiguity.
2 Related work
2.1 Optimization-based mesh estimation
Before deep learning dominates this field, 3D human mesh estimation [5, 29, 15, 30, 31] is mainly optimization-based, which optimizes the parameters of the human mesh models to match the observations. For example, Loper et al.  [5] propose MoSh that optimizes the SMPL [12] parameters to align the mesh with the 3D marker positions. It is usually used to get GT 3D meshes for benchmark datasets because of its high accuracy. Later works propose to optimize the model parameters or mesh vertices based on 2D image cues [29, 15, 30, 31, 32]. They extract intermediate representations such as 2D skeletons from the images and optimize the mesh model by minimizing the discrepancy between the model projection and the intermediate representations such as the 2D skeletons. These methods are usually sensitive to initialization and suffer from local optimum.
Figure 2: Left: The learned virtual markers (blue balls) in the back and front views. The grey balls mean they are invisible in the front view. The virtual markers act similarly to physical body markers and approximately outline the body shape. Right: Mesh estimation results by our approach VMarker, from left to right are input image, estimated 3D mesh overlayed on the image, and three different viewpoints showing the estimated 3D mesh with our intermediate predicted virtual markers (blue balls), respectively.
2.2 Learning-based mesh estimation
Recently, most studies follow the learning-based framework and achieve promising results. Existing learning-based methods can roughly be divided into two categories: deterministic mesh estimation and probabilistic mesh estimation. A considerable amount of deterministic estimation work [11, 6, 33, 34, 35, 36, 17, 37, 38, 39, 40, 10, 1] aims at estimating a single 3D human mesh from a monocular image. However, considering the inherent depth ambiguity in monocular settings, an increasing number of works propose the probabilistic modeling framework which estimates multiple feasible solutions [41, 42, 43, 44, 45] to enhance the robustness of the model.
2.2.1 Deterministic mesh estimation
Pioneer works propose to use deep networks [11, 6, 33, 36] to regress the SMPL [12] parameters from image features. However, learning the mapping from the image space to the parameter space is highly non-linear [17]. In addition, they suffer from the misalignment between the meshes and image pixels [38]. These problems make it difficult to learn an accurate yet generalizable model.
Some works propose to introduce proxy tasks to get intermediate representations first, hoping to alleviate the learning difficulty. In particular, intermediate representations of physical markers [40], IUV images [34, 38, 37, 35], body part segmentation masks [46, 47, 15, 39] and body skeletons [48, 1, 10, 24] have been proposed. In particular, THUNDR [40] first estimates the 3D locations of physical markers from images and then reconstructs the mesh from the 3D markers. The physical markers can be interpreted as a simplified representation of body shape and pose. Although it is very accurate, it cannot be applied to wild images without markers. In contrast, body skeleton is a popular human representation that can be robustly detected from wild images. Choi et al.  [1] propose to first estimate the 3D skeletons, and then estimate the intact mesh from them. However, accurate body shapes are difficult to be recovered from the oversimplified 3D skeletons.
2.2.2 Probabilistic mesh estimation
Despite the extensive exploration of multi-hypothesis 3D pose estimation [49, 50, 51, 52, 53, 54, 55], research on probabilistic 3D human mesh estimation remains relatively underdeveloped. Most recent approaches [41, 56, 42, 44, 45] predict distributions over SMPL [12] parameters conditioned on a 2D image. Biggs et al.  [41] expand upon HMR [6] to predict a categorical distribution over SMPL parameters and obtain multiple hypotheses discreetly. Sengupta et al.  [56] and ProPose [42] utilize the matrix Fisher distribution to model body pose rotation distributions. ProHMR [44] and HuManiFlow [43] employ Normalizing Flows [57, 58] to model the feasible 3D human model parameter distribution. More recently, HMDiff [45] adopts diffusion models [26] to generate multiple plausible human meshes.
Our work belongs to the learning-based class, offering both deterministic and probabilistic approaches. The deterministic approach, VMarker, introduces virtual markers as a novel intermediate representation, improving upon traditional physical markers or skeletons to better disambiguate pose and shape with more expressiveness. Our probabilistic framework, VMarker-Pro, employs diffusion models to produce diverse and plausible virtual marker estimates, from which multiple full mesh estimates that align with the image cues can be obtained.
2.3 Diffusion models
Denoising Diffusion Probabilistic Models (DDPMs) [59, 26] are a class of generative models that garner attention for their ability to generate high-quality images [26]. DDPMs employ a forward process where noise is incrementally added to the data until a noise distribution is reached, paired with a reverse process that incrementally denoises to generate data from noise. The strength of DDPMs lies in their iterative refinement, effectively capturing the target data distribution and enabling the synthesis of new samples. With the development of acceleration [26, 60] and enhancement [61, 62], the versatility of DDPMs has been demonstrated across various tasks beyond image generation [26, 60], contributing to innovations in video generation [63, 64, 65], semantic segmentation [66], motion generation [67, 68, 69], etc.
In the realm of 3D mesh estimation, existing methods [45] apply DDPMs to model a set of downsampled mesh vertices, which is hard to detect from images [17, 28]. Our approach utilizes DDPMs to model the distribution of virtual markers, an efficient and expressive intermediate representation. This reduces the complexity of the modeling, resulting in more robust and accurate estimations.
3 Method
In this section, we describe the details of our approach. First, Section 3.1 introduces how we learn the virtual marker representation from mocap data. Based on the virtual marker, we present the deterministic framework VMarker for mesh estimation from an image and its training details in Section 3.2. Then we extend VMarker to a probabilistic mesh estimation framework, VMarker-Pro in Section 3.3 and introduce its training details as well as the inference strategy.
3.1 The virtual marker representation
We represent a 3D mesh by a vector of vertex positions $\mathbf{x}\in\mathbb{R}^{3M}$ where $M$ is the number of mesh vertices. Denote a mocap dataset such as [13] with $N$ meshes as $\overset{\frown}{\mathbf{X}}=[\mathbf{x}_{1},\,...,\,\mathbf{x}_{N}]\in\mathbb{R}^{3M\times N}$. To unveil the latent structure among vertices, we reshape it to $\mathbf{X}\in\mathbb{R}^{3N\times M}$ with each column $\mathbf{x}_{i}\in\mathbb{R}^{3N}$ representing all possible positions of the $i^{\text{th}}$ vertex in the dataset [13].
The mesh $\mathbf{X}$ can be approximated by a low-rank matrix because the mesh representation is smooth and redundant where some vertices can be accurately reconstructed by others. Inspired by this, we aim to learn $K$ virtual markers from $M$ mesh vertices $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{K}]\in\mathbb{R}^{3N\times K}$, as shown in Figure 2, that try to satisfy the following two requirements to the greatest extent. First, they can accurately reconstruct the intact mesh $\mathbf{X}$ by their linear combinations: $\mathbf{X}=\mathbf{Z}\mathbf{A}$, where $\mathbf{A}\in\mathbb{R}^{K\times M}$ is a coefficient matrix that encodes the spatial relationship between the virtual markers and the mesh vertices. Second, they should have distinguishable visual patterns in images so that they can be easily detected from images. Ideally, they can be on the body surface as the meshes. While it seems natural to apply singular value decomposition (SVD) [70] to $\mathbf{X}$ to compute the eigenvectors as virtual markers for reconstructing others, there is no guarantee that the virtual markers correspond to the mesh vertices, making them difficult to be detected from images.
To that end, we apply archetypal analysis [25, 71] to learn $\mathbf{Z}$ by minimizing a reconstruction error with two additional constraints: (1) each vertex $\mathbf{x}_{i}$ can be reconstructed by convex combinations of $\mathbf{Z}$, and (2) each marker $\mathbf{z}_{i}$ should be convex combinations of the mesh vertices $\mathbf{X}$:
[…]
where $\mathbf{A}=[\bm{\alpha}_{1},...,\bm{\alpha}_{M}]\in\mathbb{R}^{K\times M}$, each $\bm{\alpha}$ resides in the simplex ${\Delta}_{K}\triangleq\{\bm{\alpha}\in\mathbb{R}^{K}\,\mathrm{s.t.}\,\bm{\alpha}\succeq 0\,\text{and}\,{||\bm{\alpha}||}_{1}=1\}$, and $\mathbf{B}=[\bm{\beta}_{1},...,\bm{\beta}_{K}]\in\mathbb{R}^{M\times K}$, $\bm{\beta}_{j}\in{\Delta}_{M}$. We adopt Active-set algorithm [71] to solve objective (Eq. 1) and obtain the learned virtual markers $\mathbf{Z}=\mathbf{X}\mathbf{B}\in\mathbb{R}^{3N\times K}$. As shown in [25, 71], the two constraints encourage the virtual markers $\mathbf{Z}$ to unveil the latent structure among vertices, therefore they learn to be close to the extreme points of the mesh and located on the body surface as much as possible.
Figure 3: Overview of our deterministic (top) and probabilistic (bottom) framework. Top: The deterministic framework VMarker. Given an input image, it first estimates the 3D virtual marker (VM) positions $\mathbf{P}$. Then we update the coefficient matrix $\mathbf{A}$ based on the estimation confidence scores $\mathbf{C}$ of the VM. Finally, the complete human mesh can be simply recovered by linear multiplication $\mathbf{M}=\mathbf{P}\mathbf{A}$. Bottom: The probabilistic framework VMarker-Pro. Building upon VMarker  we formulate the 3D VM estimation as a reverse denoising process using diffusion models. We introduce a denoiser that learns to remove the added noise from a noisy 3D VM $\hat{\mathbf{P}}(t)$ during training, conditioned on the 2D image feature $\mathbf{F}$ and the estimated 2D VMs $\mathbf{P}^{2d}$, i.e.  the first two dimensions of $\mathbf{P}$. At inference, we iteratively removes noise from random noise $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ according to Eq. 16, yielding a denoised 3D VMs $\mathbf{P}(0)$. Finally, we obtain the estimated mesh $\mathbf{M}=\mathbf{P}(0)\mathbf{A}$. VMarker-Pro can generate multiple feasible solutions aligned with 2D image cues during inference.
Post-processing. Since human body is left-right symmetric, we adjust $\mathbf{Z}$ to reflect the property. We first replace each $\mathbf{z}_{i}\in\mathbf{Z}$ by its nearest vertex on the mesh and obtain $\widetilde{\mathbf{Z}}\in\mathbb{R}^{3\times K}$. This is done on a template mesh which is in a canonical pose. This step allows us to compute the left or right counterpart of each marker. Then we replace the markers in the right body with the symmetric vertices in the left body and obtain the symmetric markers $\widetilde{\mathbf{Z}}^{sym}\in\mathbb{R}^{3\times K}$. Finally, we update $\mathbf{B}$ and $\mathbf{A}$ by minimizing $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ subject to $\widetilde{\mathbf{Z}}^{sym}=\mathbf{X}\widetilde{\mathbf{B}}^{sym}$ over all $N$ mesh data. More details are elaborated in the supplementary.
Figure 2 shows the virtual markers learned on the mocap dataset [13] after post-processing. They are similar to the physical markers and approximately outline the body shape which agrees with our expectations. They are roughly evenly distributed on the surface of the body, and some of them are located close to the body keypoints, which have distinguishable visual patterns to be accurately detected. Table I shows the reconstruction errors of using original markers $\mathbf{X}\mathbf{B}$ and the symmetric markers $\mathbf{X}\widetilde{\mathbf{B}}^{sym}$. Both can reconstruct meshes accurately.
TABLE I: The reconstruction errors using the original and the symmetric sets of markers on the H3.6M dataset [13], respectively. The errors are small indicating that they are sufficiently expressive and can reconstruct all vertices accurately.
Type Formula Reconst. Error (mm) $\downarrow$
Original $||\mathbf{X}-\mathbf{X}\mathbf{B}\mathbf{A}||^{2}_{F}$ 11.67
Symmetric $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ 10.98
3.2 Deterministic mesh estimation framework
3.2.1 VMarker
On top of the virtual markers, we present a simple yet effective framework VMarker for end-to-end 3D human mesh estimation from a single image. As shown in Figure 3 (top), it consists of two branches (yellow and orange blocks). The first branch (yellow block) uses a volumetric CNN [18] to estimate the 3D virtual marker (VM) positions $\mathbf{P}$, and the second branch (orange block) reconstructs the full mesh $\mathbf{M}$ by predicting a coefficient matrix $\mathbf{A}$:
$\mathbf{M}=\mathbf{P}\mathbf{A}.$ (2)
We will describe the two branches in more detail.
3D VM estimation. We train a neural network to estimate a 3D heatmap $\mathbf{H}=[\mathbf{H}_{1},\,...,\,\mathbf{H}_{K}]\in\mathbb{R}^{K\times D\times H\times W}$ from an image. The heatmap encodes per-voxel likelihood of each marker. There are $D\times H\times W$ voxels in total which are used to discretize the 3D space. The 3D position $\mathbf{P}_{z}\in\mathbb{R}^{3}$ of each marker is computed as the center of mass of the corresponding heatmap $\mathbf{H}_{z}$ [18] as follows:
$\mathbf{P}_{z}=\sum_{d=1}^{D}\sum_{h=1}^{H}\sum_{w=1}^{W}(d,h,w)\cdot\mathbf{H}_{z}(d,h,w).$ (3)
The positions of all markers are represented as $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\cdots,\mathbf{P}_{K}]$.
3D mesh estimation. Ideally, if we have accurate estimates for all virtual markers $\mathbf{P}$, then we can recover the complete mesh by simply multiplying $\mathbf{P}$ with a fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ with sufficient accuracy as validated in Table I. However, in practice, some markers may have large estimation errors because they may be occluded in the monocular setting. Note that this happens frequently. For example, the markers in the back will be occluded when a person is facing the camera. As a result, inaccurate markers positions may bring large errors to the final mesh if we directly multiply them with the fixed matrix $\widetilde{\mathbf{A}}^{sym}$.
Our solution is to rely more on those accurately detected markers. To that end, we propose to update the coefficient matrix based on the estimation confidence scores $\mathbf{C}$ of the markers. In practice, we simply take the heatmap score at the estimated positions of each marker obtained by interpolation. For convenience, we denote this process as $\mathbf{C}_{z}=\mathbf{H}_{z}(\mathbf{P}_{z})$. We feed the confidence scores to an updating network which is a single fully-connected (FC) layer to obtain the coefficient matrix $\mathbf{A}$. Then the mesh is reconstructed by linear multiplication $\mathbf{M}=\mathbf{P}\mathbf{A}$.
3.2.2 Training
We train the whole network end-to-end in a supervised way. The overall loss function is defined as:
$\mathcal{L}_{vm}=\lambda_{3d}\mathcal{L}_{3d}+\lambda_{c}\mathcal{L}_{conf}+\lambda_{m}\mathcal{L}_{mesh}.$ (4)
Virtual marker loss. We define $\mathcal{L}_{3d}$ as the $L_{1}$ distance between the predicted 3D virtual markers $\mathbf{P}$ and the GT $\hat{\mathbf{P}}$ as follows:
$\mathcal{L}_{3d}=\|\mathbf{P}-\hat{\mathbf{P}}\|_{1}.$ (5)
Note that it is easy to get GT markers $\hat{\mathbf{P}}$ from GT meshes as stated in Section 3.1 without additional manual annotations.
Confidence loss. We also require that the 3D heatmaps have reasonable shapes, therefore, the heatmap score at the voxel containing the GT marker position $\hat{\mathbf{P}}_{z}$ should have the maximum value as in the previous work [72]:
$\mathcal{L}_{conf}=-\sum_{z=1}^{K}log(\mathbf{H}_{z}(\hat{\mathbf{P}}_{z})).$ (6)
Mesh loss. Following [17], we define $\mathcal{L}_{mesh}$ as a weighted sum of four losses:
$\mathcal{L}_{mesh}=\mathcal{L}_{vertex}+\mathcal{L}_{pose}+\mathcal{L}_{normal}+\lambda_{e}\mathcal{L}_{edge}.$ (7)
  • Vertex coordinate loss. We adopt $L_{1}$ loss between predicted 3D mesh coordinates $\mathbf{M}$ with GT mesh $\hat{\mathbf{M}}$ as:
    $\mathcal{L}_{vertex}=\|\mathbf{M}-\hat{\mathbf{M}}\|_{1}.$ (8)
  • Pose loss. We use $L_{1}$ loss between the 3D landmark joints regressed from mesh $\mathbf{M}\mathcal{J}$ and the GT joints $\hat{\mathbf{M}}\mathcal{J}$ as:
    $\mathcal{L}_{pose}=\|\mathbf{M}\mathcal{J}-\hat{\mathbf{M}}\mathcal{J}\|_{1},$ (9)
    where $\mathcal{J}\in\mathbb{R}^{M\times J}$ is a pre-defined joint regression matrix in SMPL model [29] and $J$ is the number of joints.
  • Surface losses. To improve surface smoothness [73], we supervise the normal vector of a triangle face with GT normal vectors by $\mathcal{L}_{normal}$ and the edge length of the predicted mesh with GT length by $\mathcal{L}_{edge}$:
    $\mathcal{L}_{normal}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\left<\frac{\mathbf{M}_{i}-\mathbf{M}_{j}}{\|\mathbf{M}_{i}-\mathbf{M}_{j}\|_{2}},\hat{\mathbf{n}}_{f}\right>\right|,$ (10)
    $\mathcal{L}_{edge}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\|\mathbf{M}_{i}-\mathbf{M}_{j}\|_{2}-\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}\right|.$
    where $f$ and $\hat{\mathbf{n}}_{f}$ denote a triangle face in the mesh and its GT unit normal vector, respectively. $\mathbf{M}_{i}$ denote the $i^{th}$ vertex of $\mathbf{M}$. $\hat{}$ denotes GT.
3.3 Probabilistic mesh estimation framework
3.3.1 VMarker-Pro
As shown in Figure 2, virtual markers are often self-occluded, thus there may exist multiple viable solutions. An incorrect estimation may introduce errors into the final mesh estimate. Therefore, we propose a probabilistic framework for modeling the estimation of 3D VMs, as illustrated in Figure 3 (bottom). By leveraging the powerful modeling capabilities of diffusion models [26, 74] on data distributions, we generate multiple feasible 3D VM estimates given an image condition. Specifically, we model the estimation of 3D VMs as a conditional reverse denoising process that starts with Gaussian noise and progressively denoises it to obtain the 3D VM using a denoiser network. In the following, we detail the problem formulation of the 3D VM estimation using diffusion models, including the definition of the forward diffusion process and the reverse denoising process. Then, we introduce the architecture design of the denoiser network.
Problem formulation. We formulate the 3D VM estimation as a reverse denoising process [26], as shown in the inference workflow at the bottom of Figure 3. Starting with a Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$, we design a denoiser network to iteratively estimate the clean 3D VM. To achieve this, the framework includes two core processes: a forward diffusion process and a reverse denoising process. (1) The forward diffusion process disturbs a clean 3D VM, i.e.  the ground-truth (GT) 3D VM $\hat{\mathbf{P}}$ from the data distribution, by incrementally adding noise to generate a series of noisy 3D VMs and ultimately reaches a Gaussian prior. (2) Then the reverse denoising process iteratively denoises the Gaussian noise into clean data in a step-by-step manner. To guide the denoising process, we condition it on the 2D image cues $\mathbf{F}$ and the estimated 2D VMs $\mathbf{P}^{2d}\in\mathbb{R}^{2\times K}$, which is the first two dimensions of the estimated 3D VMs $\mathbf{P}$ by VMarker. By formulating the estimation as a reverse denoising process, we leverage the generative capabilities of diffusion models, framing the task as a probabilistic estimation naturally.
Forward diffusion process. Starting from a clean 3D VM $\hat{\mathbf{P}}$ drawn from the GT data distribution, also denoted as $\hat{\mathbf{P}}(0)$, we construct a time-dependent diffusion process by incrementally adding noise to $\hat{\mathbf{P}}(0)$ to generate a series of noisy 3D VMs $\{\hat{\mathbf{P}}(t)\}_{t=0}^{T}$, indexed by a discrete timestep $t\in\{0,1,...,T\}$, where $T$ denotes the total number of timesteps. Formally, we define a noisy 3D VM $\hat{\mathbf{P}}(t)$ at timestep $t$ as:
$q(\hat{\mathbf{P}}(1:T)|\hat{\mathbf{P}}(0))$ $:=\prod_{t=1}^{T}q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(t-1)),$ (11)
$q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(t-1))$ $:=\mathcal{N}(\hat{\mathbf{P}}(t);\sqrt{1-\beta_{t}}\hat{\mathbf{P}}(t-1),\beta_{t}\mathbf{I}),$ (12)
where $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule [26]. Leveraging the additive property of independent Gaussian distributions and reparameterization techniques [75], we can obtain the perturbation $\hat{\mathbf{P}}(t)$ from a clean 3D VM $\hat{\mathbf{P}}(0)$ directly:
$q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(0)):=\sqrt{\overline{\alpha}_{t}}\hat{\mathbf{P}}(0)+\sqrt{1-\overline{\alpha}_{t}}\hat{\bm{\epsilon}},$ (13)
where $\alpha_{t}:=1-\beta_{t}$, $\overline{\alpha_{t}}:=\prod_{u=1}^{t}\alpha_{u}$, and $\hat{\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ as a Gaussian noise. Following previous work [26], we disturb any clean 3D VM $\hat{\mathbf{P}}(0)$ to a Gaussian distribution $\hat{\mathbf{P}}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
Reverse denoising process. Starting from a Gaussian distribution, denoted as $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, we reverse the diffusion process by incrementally denoising it to get a clean 3D VM $\mathbf{P}(0)$. We adopt Denoising Diffusion Implicit Models (DDIM) [60] to accelerate the reverse denoising process with fewer timesteps while keeping the denoising quality. Formally, following DDIM [60], we define a timestep subset $\tau\subset\{0,1,...,T\}$, and the reverse process is defined as:
$\begin{split}\mathbf{P}({\tau_{i-1}})&=\sqrt{\overline{\alpha}_{\tau_{i-1}}}(\frac{\mathbf{P}({\tau_{i}})-\sqrt{1-\overline{\alpha}_{\tau_{i}}}\bm{\epsilon}_{\tau_{i}}}{\sqrt{\overline{\alpha}_{\tau_{i}}}})\\ &+\sqrt{1-\overline{\alpha}_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\bm{\epsilon}_{\tau_{i}}+\sigma_{\tau_{i}}\hat{\bm{\epsilon}}_{\tau_{i}},\end{split}$ (14)
where $\tau_{i},\tau_{i-1}$ are the adjacent timesteps in the subset $\tau$. $\sigma_{\tau_{i}}(\eta)=\eta\sqrt{(1-\overline{\alpha}_{\tau_{i-1}})/({1-\overline{\alpha}_{\tau_{i}}})}\sqrt{1-\overline{\alpha}_{\tau_{i}}/\overline{\alpha}_{\tau_{i-1}}}$, where $\eta\in\mathbb{R}_{\geq 0}$ is a hyperparameter to control the generation process. $\hat{\bm{\epsilon}}_{\tau_{i}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is standard Gaussian noise.
In order to simulate Eq. 14 to perform the denoising process, we need to know $\bm{\epsilon}_{\tau_{i}}$ for all timestep $\tau_{i}$. Therefore, we train a time-dependent denoiser network to estimate it as shown at the bottom block of Figure 3. Instead of directly predicting $\bm{\epsilon}_{\tau_{i}}$ as the original formulation by Ho et al.  [26], following previous work [76, 68], we predict the clean signal itself, i.e.  $\mathbf{P}(0)$, and obtain the noise $\bm{\epsilon}_{\tau_{i}}$ through:
$\bm{\epsilon}_{\tau_{i}}=\frac{\mathbf{P}({\tau_{i}})-\sqrt{\overline{\alpha}_{{\tau_{i}}}}\mathbf{P}(0)}{\sqrt{1-\overline{\alpha}_{\tau_{i}}}}.$ (15)
By combining Eq. 15 and Eq. 14, we can get the denoised $\mathbf{P}({\tau_{i-1}})$ as follows:
$\mathbf{P}({\tau_{i-1}})$ $=\sqrt{\overline{\alpha}_{\tau_{i-1}}}\mathbf{P}(0)$ (16)
$+\sqrt{1-\overline{\alpha}_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\frac{\mathbf{P}({\tau_{i}})-\sqrt{\overline{\alpha}_{{\tau_{i}}}}\mathbf{P}(0)}{\sqrt{1-\overline{\alpha}_{\tau_{i}}}}$
$+\sigma_{\tau_{i}}\hat{\bm{\epsilon}}_{\tau_{i}}.$
Denoiser architecture design. Figure 4 presents an overview of the time-dependent denoiser network architecture. The denoiser aims to estimate the clean 3D VM $\mathbf{P}(0)$ given a noisy 3D VM $\hat{\mathbf{P}}(t)$ at timestep $t$ as input. To guide the prediction, we concatenate $\hat{\mathbf{P}}(t)$ with the estimated 2D VM $\mathbf{P}^{2d}$, and then transform them to feature $\mathbf{F}^{vm}\in\mathbb{R}^{K\times C}$ by a multilayer perceptron (MLP), where $C$ denotes the number of feature channel. We then deploy a network with $B$ blocks to process the VM features $\mathbf{F}^{vm}$. Each block is composed of two Locally Connected Network (LCN) units [77, 78] sequentially. Within each unit, an LCN layer [77, 78] is followed by Group Normalization (GN) [79], a SiLU activation function [80], and Dropout [81] for regularization. The LCN layer, as proposed in [77, 78], is akin to an enhanced version of MLP, offering more effective feature processing across virtual markers.
To improve the alignment of the denoised 3D VM with 2D image cues, we extract 2D image features $\mathbf{F}\in\mathbb{R}^{H_{b}\times W_{b}\times C_{b}}$ using the CNN backbone [18]. These features are sampled according to the estimated 2D VMs $\mathbf{P}^{2d}$ and subsequently transformed into a new feature space $\mathbf{F}^{2d}\in\mathbb{R}^{K\times C}$ through a linear layer, as illustrated in Figure 4. We then fuse $\mathbf{F}^{2d}$ into the network by Cross-Attention [82]. To achieve this, we slightly modify the LCN layer by incorporating an attention module [82].
In addition, we process the current timestep $t$ using an MLP to be $\mathbf{F}^{t}\in\mathbb{R}^{1\times C}$ and add it to the processed VM features $\mathbf{F}^{vm}$ to make the denoiser time-dependent following common practice [74, 53]. Finally, an FC layer is deployed to estimate the clean signal $\mathbf{P}(0)$. Please refer to the supplementary for more details.
3.3.2 Training
The whole framework is trained in an end-to-end manner. The overall loss function is defined as:
$\mathcal{L}_{vmpro}=\mathcal{L}_{vm}+\lambda_{diffusion}\mathcal{L}_{diffusion},$ (17)
where $\mathcal{L}_{vm}$ is the loss function for VMarker, as defined in Eq. 4. Following common practice in training diffusion models [26, 68], we randomly sample a timestep $t\in\{0,1,...,T\}$ to get the noisy 3D VM $\hat{\mathbf{P}}(t)$ according to Eq. 13 and deploy $\mathcal{L}_{diffusion}$ as the $L_{1}$ distance between the predicted clean 3D VM $\mathbf{P}(0)$ and the GT $\hat{\mathbf{P}}(0)$:
$\mathcal{L}_{diffusion}=\mathbb{E}_{t,\hat{\mathbf{P}}(0),\hat{\bm{\epsilon}}}[\|\mathbf{P}(0)-\hat{\mathbf{P}}(0)\|_{1}].$ (18)
3.3.3 Inference
As shown in Figure 3, at inference, the denoiser iteratively removes noise from a Gaussian noise $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ using DDIM [60] according to Eq. 16, yielding a denoised 3D VM $\mathbf{P}(0)$. Finally, we obtain a mesh estimate $\mathbf{M}=\mathbf{P}(0)\mathbf{A}$. By sampling $S$ Gaussian noises, we can obtain $S$ feasible mesh estimates.
Figure 4: Architecture of the denoiser network, which predicts the clean 3D VM $\mathbf{P}(0)$ from the noisy 3D VM $\hat{\mathbf{P}}(t)$. To guide the direction of denoising, we additionally input 2D features $\mathbf{F}$, the current timestep $t$, and the 2D VM $\mathbf{P}^{2d}$ predicted by VMarker. We concatenate $\mathbf{P}^{2d}$ with the noisy 3D VM $\hat{\mathbf{P}}(t)$ to facilitate training and fuse the sampled 2D features $\mathbf{F}$ through a Cross-Attention mechanism [82]. We utilize a network comprising $B$ blocks to process the features, where each block contains two Locally Connected Network (LCN) units [77, 78]. Each LCN units consists of one LCN layer [77, 78], followed by Group Normalization (GN) [79], SiLU activation [80], and Dropout [81]. Finally, an FC layer is used to predict the clean 3D VM $\mathbf{P}(0)$.
4 Experiments
TABLE II: Comparison to the deterministic state-of-the-arts on H3.6M [13] and 3DPW [14] datasets. means using temporal cues. The methods are not strictly comparable because they may have different backbones and training datasets. We provide the numbers only to show proof-of-concept results.
Method Venue Intermediate H3.6M [13] 3DPW [14]
[…]
HMMR [83] CVPR’19 - - - 56.9 139.3 116.5 72.6
DSD-SATN [48] ICCV’19 3D skeleton - 59.1 42.4 - - 69.5
VIBE [84] CVPR’20 - - 65.9 41.5 99.1 82.9 51.9
TCMR [85] CVPR’21 - - 62.3 41.1 102.9 86.5 52.7
MAED [24] ICCV’21 3D skeleton - 56.3 38.7 92.6 79.1 45.7
SMPLify [29] ECCV’16 2D skeleton - - 82.3 - - -
HMR [6] CVPR’18 - 96.1 88.0 56.8 152.7 130.0 81.3
GraphCMR [16] CVPR’19 3D vertices - - 50.1 - - 70.2
SPIN [33] ICCV’19 - - - 41.1 116.4 96.9 59.2
DenseRac [34] ICCV’19 IUV image - 76.8 48.0 - - -
DecoMR [38] CVPR’20 IUV image - 60.6 39.3 - - -
ExPose [86] ECCV’20 - - - - - 93.4 60.7
Pose2Mesh [1] ECCV’20 3D skeleton 85.3 64.9 46.3 106.3 88.9 58.3
I2L-MeshNet [17] ECCV’20 3D vertices 65.1 55.7 41.1 110.1 93.2 57.7
PC-HMR [36] AAAI’21 3D skeleton - - - 108.6 87.8 66.9
HybrIK [10] CVPR’21 3D skeleton 65.7 54.4 34.5 86.5 74.1 45.0
METRO [87] CVPR’21 3D vertices - 54.0 36.7 88.2 77.1 47.9
ROMP [88] ICCV’21 - - - - 108.3 91.3 54.9
Mesh Graphormer[89] ICCV’21 3D vertices - 51.2 34.5 87.7 74.7 45.6
PARE [47] ICCV’21 Segmentation - - - 88.6 74.5 46.5
THUNDR [40] ICCV’21 3D markers - 55.0 39.8 88.0 74.8 51.5
PyMaf [37] ICCV’21 IUV image - 57.7 40.5 110.1 92.8 58.9
OCHMR [90] CVPR’22 2D heatmap - - - 107.1 89.7 58.3
3DCrowdNet [91] CVPR’22 3D skeleton - - - 98.3 81.7 51.5
CLIFF [92] ECCV’22 - - 47.1 32.7 81.2 69.0 43.0
FastMETRO [93] ECCV’22 3D vertices - 52.2 33.7 84.1 73.5 44.6
VisDB [94] ECCV’22 3D vertices - 51.0 34.5 85.5 73.5 44.9
VMarker (Ours) [28] CVPR’23 Virtual marker 58.0 47.3 32.0 77.9 67.5 41.3
TABLE III: Comparison to the probabilistic state-of-the-arts on H3.6M [13] and 3DPW [14] datasets. The methods are not strictly comparable because they may have different backbones and training datasets. We provide the numbers only to show proof-of-concept results.
Method Hypothesis H3.6M [13] 3DPW [14]
[…]
Biggs et al.  [41] NeurIPS’20 10 - 59.2 42.2 - 79.4 56.6
25 - 58.2 42.2 - 75.8 55.6
Sengupta et al.  [56] ICCV’21 25 - - - - 75.1 47.0
ProHMR [44] ICCV’21 10 - - - - 88.9 55.0
25 - - - - 85.1 52.1
100 - - - - 80.1 48.1
200 - - - - 77.9 46.5
HuManiFlow [43] CVPR’23 10 - - - - 75.6 47.9
25 - - - - 71.9 44.5
100 - - - - 65.1 39.9
200 - - - - 64.5 38.8
HMDiff [45] ICCV’23 25 - 49.3 32.4 82.4 72.7 44.5
10 55.8 45.4 31.3 77.3 66.6 42.5
25 52.9 42.9 29.8 72.9 63.2 40.0
100 49.2 39.9 27.6 67.4 58.2 36.4
VMarker-Pro (Ours) 200 47.7 38.6 26.6 64.7 56.2 35.0
4.1 Datasets and metrics
H3.6M [13]. We use (S1, S5, S6, S7, S8) for training and (S9, S11) for testing. For the deterministic estimation task, as in [6, 1, 87, 89], we report Mean Per Joint Position Error (MPJPE) and PA-MPJPE for poses that are derived from the estimated meshes. We also report Mean Per Vertex Error (MPVE) for the whole mesh. For the probabilistic estimation task, we report the minimum errors over $S$ hypotheses to measure the estimated distribution accuracy following a standard practice in recent works [43, 45].
3DPW [14] is collected in natural scenes. Following previous works [87, 89, 47, 40, 43, 45], we use the train set of 3DPW to learn the model and evaluate on the test set. The same evaluation metrics as H3.6M are used. We further evaluate the performance of our probabilistic framework VMarker-Pro on two subsets of 3DPW, i.e.  3DPW-OC [14, 95] and 3DPW-PC [14, 88] which are composed by object- and person-specific occlusion, respectively. To ensure a fair comparison, as suggested by [96], we do not use the 3DPW training set in these particular evaluations.
SURREAL [27] is a large-scale synthetic dataset with GT SMPL annotations and has diverse samples in terms of body shapes, backgrounds, etc.  We use its training set to train a model and evaluate the test split following [1]. The same evaluation metrics as H3.6M [13] are reported.
4.2 Implementation Details
We learn $64$ virtual markers on the H3.6M [13] training set. We use the same set of markers for all datasets instead of learning a separate set for each one. Following [6, 1, 17, 40, 16, 84, 89, 87], we conduct mix-training by using MPI-INF-3DHP [97], UP-3D [15], and COCO [98] training set for experiments on the H3.6M and 3DPW datasets.
We adapt a 3D pose estimator [18] with HRNet-W48 [99] as the image feature backbone for estimating the 3D virtual markers. We set the number of voxels in each dimension to be $64$, i.e.  $D=H=W=64$ for 3D heatmaps. Following [6, 16, 17], we crop every single human region from the input image and resize it to $256\times 256$. The dimensions of the 2D feature map are $64$, i.e.  $H_{b}=W_{b}=64$. The 2D feature map has $C_{b}=48$ feature channels. The processed $\mathbf{F}^{2d}$ feature channels are set to be $C=64$. The denoiser network has $B=3$ blocks. To train VMarker, we use Adam [100] optimizer to train the whole framework for $40$ epochs with a batch size of $32$. The learning rates for the 3D VM estimation branch and the updating branch are set to $5\times 10^{-4}$ and $1\times 10^{-3}$, respectively, which are decreased by half after the $30^{th}$ epoch. To train VMarker-Pro, we use Adam [100] optimizer to train the whole framework for $50$ epochs with a batch size of $80$. The learning rates for the 3D VM estimation branch, the updating branch, and the denoiser are set to $4\times 10^{-5}$, $2\times 10^{-4}$, and $2\times 10^{-3}$, respectively. They are decayed by $0.5$, $0.5$, and $0.1$ after $30^{th}$ and $40^{th}$ epochs, respectively. For inference, we employ DDIM [60] strategy and generate each hypothesis in $T=10$ steps with $\eta=0$. Please check the supplementary for more details.
TABLE IV: Comparison to the state-of-the-arts on SURREAL [27] dataset. means training on the test split with 2D supervisions. “Skel. + Seg.” means using skeleton and segmentation together. $S$ is the number of hypotheses.
[…]
Representation
HMR [6] - - 85.1 73.6 55.4
BodyNet [46] Skel. + Seg. - 65.8 - -
GraphCMR [16] 3D vertices - 103.2 87.4 63.2
SPIN [33] - - 82.3 66.7 43.7
DecoMR [38] IUV image - 68.9 52.0 43.0
Pose2Mesh [1] 3D skeleton - 68.8 56.6 39.6
PC-HMR [36] 3D skeleton - 59.8 51.7 37.9
DynaBOA [101] - - 70.7 55.2 34.0
VMarker (Ours) Virtual marker - 44.7 36.9 28.9
10 44.3 36.6 28.0
VMarker-Pro 25 42.0 34.8 26.2
(Ours) 100 38.1 31.0 23.5
Virtual marker 200 36.5 29.8 22.4
TABLE V: Comparison to the state-of-the-art methods on the challenging occlusion-specific 3DPW-OC [14, 95] and 3DPW-PC [14, 88] datasets. The top and bottom blocks are deterministic and probabilistic methods, respectively.
Method Venue Hypothesis 3DPW-OC [14, 95] 3DPW-PC [14, 88]
[…]
SPIN [33] ICCV’19 - 121.4 95.5 60.7 159.8 122.1 77.4
I2L-MeshNet [17] ECCV’20 - 129.5 92.0 61.4 160.2 117.3 80.0
PyMAF [37] ICCV’21 - 113.7 89.6 59.1 154.6 117.5 74.5
ROMP [88] ICCV’21 - - 91.0 62.0 152.8 117.9 79.7
PARE [47] ICCV’21 - 101.5 83.5 57.0 122.4 96.8 64.5
OCHMR [90] CVPR’22 - 145.9 112.2 75.2 149.6 117.5 77.1
3DCrowdNet [91] CVPR’22 - 101.5 83.5 57.1 114.8 90.9 64.4
JOTR [96] ICCV’23 - 92.6 75.7 52.2 109.7 86.5 58.3
VMarker (Ours) [28] CVPR’23 - 92.8 78.4 49.0 110.8 93.9 66.0
HMDiff [45] ICCV’23 25 - - - 143.1 114.2 73.5
10 90.4 77.3 48.6 109.4 89.8 60.2
25 85.4 73.3 46.0 102.9 84.7 56.6
100 77.7 67.1 42.5 94.0 78.0 52.0
VMarker-Pro (Ours) - 200 75.2 64.9 40.9 90.3 74.6 49.7
4.3 Comparison to the State-of-the-arts
4.3.1 Deterministic Track
Results on H3.6M [13]. Table II compares our approach to the deterministic state-of-the-art (SOTA) methods on the H3.6M dataset. Our method achieves superior performance. In particular, it outperforms the methods that use skeletons (Pose2Mesh [1], DSD-SATN [48]), body markers (THUNDR) [40], or IUV image [38, 37] as proxy representations, demonstrating the effectiveness of the virtual marker representation.
Results on 3DPW [14]. We compare our method to the deterministic SOTA methods on the 3DPW dataset in Table II. Our approach achieves SOTA results among all the methods, validating the advantages of the virtual marker representation over the skeleton representation used in Pose2Mesh [1], DSD-SATN [48], and other representations like IUV image used in PyMAF [37]. In particular, our approach outperforms I2L-MeshNet [17], METRO [87], and Mesh Graphormer [89] by a notable margin, suggesting that virtual markers are more suitable and effective representations than using all vertices as most of them are not discriminative enough to be accurately detected.
Results on SURREAL [27]. This dataset has more diverse samples in terms of body shapes. The results are shown in Table IV. Our approach VMarker, outperforms the SOTA methods by a notable margin, especially in terms of MPVE. Figure 1 (top) shows some challenging cases without cherry-picking. The skeleton representation loses the body shape information so the method [1] can only recover mean shapes. In contrast, our approach generates much more accurate mesh estimation results.
4.3.2 Probabilistic Track
Results on H3.6M [13] and 3DPW [14]. Table III presents a comparison between VMarker-Pro and other SOTA probabilistic methods on the H3.6M [13] and 3DPW [14] datasets. Our method significantly surpasses the existing SOTA approaches in terms of accuracy when estimating data distributions. This serves as a validation of the effectiveness of our probabilistic framework, which benefits from the generation capability of diffusion models. Furthermore, as the number of hypotheses ($S$) increases, our method consistently improves, highlighting the diverse and accurate nature of our approach when modeling data distributions. Notably, compared to HMDiff [45], which is also a diffusion model-based method, our incorporation of 2D image cues and the virtual marker representations contributes to the observed notable improvement.
Results on 3DPW-OC [14, 95] and 3DPW-PC [14, 88]. Table V shows the comparison results of our method against existing methods on the 3DPW-OC and 3DPW-PC datasets. Both 3DPW-OC [14, 95] and 3DPW-PC [14, 88] are subsets of the 3DPW [14] dataset, containing scenarios with occlusions by objects and people, respectively. It can be seen that our deterministic estimation results are comparable to the SOTA performance. When adopting the proposed probabilistic modeling approach, the accuracy of our method in estimating the data distribution is significantly improved, thanks to the enhanced accuracy in 3D virtual marker estimation. Figure 5 presents a comparative visualization between VMarker and VMarker-Pro. It can be seen that the deterministic approach, VMarker, may encounter difficulties under occlusions, e.g.  the wrongly estimated right arm and the incorrect arm length. Conversely, the probabilistic method VMarker-Pro, demonstrates a capability to accurately model realistic human poses, offering feasible estimations for the occluded parts.
Figure 5: Qualitative comparison of VMarker and VMarker-Pro on 3DPW-OC [14, 88] testset. For each mesh, we show a projecting view and a side view (with shadow). VMarker wrongly estimates the right arm when heavy occlusion happens while VMarker-Pro provides two reasonable estimates.
Results on SURREAL [27]. Table IV displays the results of our probabilistic approach on the SURREAL [27] dataset. It can be seen that when $S=10$, our probabilistic framework already surpasses existing approaches. This indicates that our method can accurately model the more diverse data distribution. As the number of hypotheses ($S$) increases, the performance of our model continues to improve.
4.4 Ablation Study on VMarker
Virtual marker representation. We compare our method to two baselines in Table VI. First, in baseline (a), we replace the virtual markers of our method with the skeleton representation. The rest are kept the same as ours (c). Our method achieves a much lower MPVE than the baseline (a), demonstrating that the virtual markers help to estimate body shapes more accurately than the skeletons. In baseline (b), we randomly sample $64$ from the $6890$ mesh vertices as virtual markers. We repeat the experiment five times and report the average number. We can see that the result is worse than ours, which is because the randomly selected vertices may not be expressive enough to reconstruct the other vertices or can not be accurately detected from images as they lack distinguishable visual patterns. The results validate the effectiveness of our learning strategy.
TABLE VI: Ablation study of the virtual marker representation for our approach VMarker on H3.6M [13] and SURREAL [27] datasets. “Skeleton” means the sparse landmark joint representation is used. “Rand virtual marker” means the virtual markers are randomly selected from all the vertices without learning. (c) is our method, where the learned virtual markers are used. MPVE error is reported.
[…]
Representation H3.6M SURREAL
(a) Skeleton 64.4 53.6
(b) Rand virtual marker 63.0 50.1
(c) Virtual marker 58.0 44.7
Figure 6: Mesh estimation results of different methods on H3.6M [13] test set. Our method VMarker with virtual marker representation gets better shape estimation results than Pose2Mesh [1] which uses skeleton representation. Note the waistline of the body and the thickness of the arm.
Figure 1 (top) shows some qualitative results on the SURREAL test set. The meshes estimated by the baseline which uses skeleton representation, i.e.  Pose2Mesh [1], have inaccurate body shapes. This is reasonable because the skeleton is oversimplified and has very limited capability to recover shapes. Instead, it implicitly learns a mean shape for the whole training dataset. In contrast, the mesh estimated by using virtual markers has much better quality due to its strong representation power and therefore can handle different body shapes elegantly. Figure 6 also shows some qualitative results on the H3.6M test set. For clarity, we also draw the intermediate representation (blue balls) in it.
Number of virtual markers. We evaluate how the number of virtual markers affects estimation quality on H3.6M [13] dataset. Figure 2 visualizes the learned 64 markers. Figure 7 visualizes the learned virtual markers when $K=16,32,96$, which are all located on the body surface and close to the extreme points of the mesh. This is expected as mentioned in Section 3.1. Table VII (GT) shows the mesh reconstruction results when we have GT 3D positions of the virtual markers in objective (Eq. 1). When we increase the number of virtual markers, both mesh reconstruction error, i.e.  MPVE, and the regressed landmark joint error, i.e.  MPJPE, steadily decrease. This is expected because using more virtual markers improves the representation power. However, using more virtual markers cannot guarantee smaller estimation errors when we need to estimate the virtual marker positions from images as in our method. This is because the additional virtual markers may have large estimation errors which affect the mesh estimation result. The results are shown in Table VII (Est). Increasing the number of virtual markers $K$ steadily reduces the MPVE errors when $K$ is smaller than $96$. However, if we keep increasing $K$, the error begins to increase. This is mainly because some of the newly introduced virtual markers are difficult to detect from images and therefore bring errors to mesh estimation.
TABLE VII: Ablation study of the different number of virtual markers ($K$) on H3.6M [13] dataset. (GT) Mesh reconstruction results when GT 3D positions of the virtual markers are used in objective (1). (Est) Mesh estimation results obtained by our method VMarker when we use different numbers of virtual markers ($K$).
$K$ GT Est
[…]
16 46.8 39.8 58.7 47.8
32 20.1 14.2 58.2 48.3
64 11.0 7.5 58.0 47.3
96 9.9 5.6 59.6 48.2
Figure 7: Visualization of the learned virtual markers when $K=16,32,96$, from left to right, respectively.
Updating coefficient matrix. We compare our method to a baseline which uses the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$. We show the quality comparison in Figure 8. We can see that the estimated mesh by (a) a fixed coefficient matrix has mostly correct pose and shape but there are some artifacts on the mesh while using the (b) updated coefficient matrix can get better mesh estimation results. As shown in Table VIII, using a fixed coefficient matrix gets larger MPVE and MPJPE errors than using the updated one. This is caused by the estimation errors of virtual markers when occlusion happens, which is inevitable since the virtual markers on the back will be self-occluded by the front body. As a result, inaccurate marker positions would bring large errors to the final mesh estimates if we directly use the fixed matrix.
TABLE VIII: Ablation study of the coefficient matrix for VMarker on H3.6M [13] dataset. “fixed” means using the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ to reconstruct the mesh.
[…]
(a) VMarker (fixed) 64.7 51.6
(b) VMarker 58.0 47.3
Figure 8: Mesh estimation comparison results when using (a) fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$, and (b) updated $\mathbf{A}$ in our method VMarker. Please zoom in to better see the details.
TABLE IX: Ablation study of the architecture design for the probabilistic method VMarker-Pro on 3DPW [14] dataset. (a-b) ablates the use of 2D feature $\mathbf{F}$ and 2D VMs $\mathbf{P}^{2d}$ in denoiser network, respectively. Ablation (c) removes the 2D VM $\mathbf{P}^{2d}$ sampling in 2D feature $\mathbf{F}$ and directly transforms the global 2D feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ for the denoiser. In ablation (d), the denoiser directly estimates the noise $\bm{\epsilon}$ following the original formula [26], while keeping the rest design the same. (e) is the proposed VMarker-Pro framework.
[…]
$\mathbf{F}$ $\mathbf{P}^{2d}$ $\mathbf{H}^{2d}$
(a) w/o $\mathbf{F}$ 80.1 68.2
(b) w/o $\mathbf{P}^{2d}$ 80.6 68.8
(c) w/o sampling in $\mathbf{F}$ 80.4 69.2
(d) $\rightarrow\bm{\epsilon}$ 80.6 68.9
(e) VMarker-Pro 76.9 66.1
Figure 9: Mesh estimation comparison results of different ablated methods of VMarker-Pro. (a) We ablate the usage of 2D image features $\mathbf{F}$. (b) We ablate the usage of 2D estimated VM $\mathbf{P}^{2d}$. (c) We remove the 2D VM $\mathbf{P}^{2d}$ sampling in 2D feature $\mathbf{F}$ and directly transform the global 2D feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ for the denoiser. (d) We change the estimation target of the denoiser to the noise $\bm{\epsilon}$ instead of the signal itself. (e) is our proposed VMarker-Pro method. We show the GT mesh and its side view on the far right. For baselines (a-e), we show each hypothesis that is denoised from a zero noise for a fair comparison. Each hypothesis presents a projecting view and a side view.
4.5 Ablation Study on VMarker-Pro
Architecture design of the denoiser. Table IX compares VMarker-Pro to four baselines. We report the metric when predicting $S=10$ hypotheses. In baseline (a) and (b), we remove the 2D feature $\mathbf{F}$ and 2D estimated VMs $\mathbf{P}^{2d}$ input for the denoiser, respectively. It can be seen that the errors increase a lot compared to the full model, which validates the effectiveness of the two guidance. In baseline (c), we remove the local sampling by 2D VMs $\mathbf{P}^{2d}$ on 2D feature $\mathbf{F}$, i.e.  the purple arrow in Figure 4. Instead, we directly transform the global 2D image feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ when feeding into the denoiser and the performance degrades notably. In ablation study (d), we alter the denoiser to estimate the noise $\bm{\epsilon}$ rather than the clean 3D VMs $\mathbf{P}(0)$, adjusting the inference process to align with the original denoising formulation by Ho et al.  [26]. This modification results in a decrease in performance. A similar observation was made in previous research by Tevet et al.  [68], where it is suggested that noise estimation might require substantially larger batch sizes to be effective. Unfortunately, we lack the resources to confirm this hypothesis in our settings.
Figure 9 presents the qualitative comparison of different designs. As observed in the image, the left arm of the subject is nearly fully extended. However, due to depth ambiguity, the estimations from baselines (a-d) align with the 2D perspective, yet inaccuracies arise in the 3D estimation of the left forearm when viewed from the side view. Our method (e), which integrates 2D local features, more accurately captures precise 3D poses, achieving estimations that most closely match the GT. This highlights the significance of our design, which not only utilizes 2D local features $\mathbf{F}$ but also leverages 2D VMs $\mathbf{P}^{2d}$ to locally sample these features effectively.
TABLE X: Results of using different denoising steps $T$ for VMarker-Pro on H3.6M [13] and 3DPW [14] datasets.
[…]
1 386.5 370.4 372.2 358.9 339.6
2 77.0 66.2 104.7 92.6 326.6
4 58.8 48.3 80.0 69.2 301.0
5 57.1 46.9 78.3 67.6 292.6
10 55.8 45.4 77.3 66.6 247.4
15 55.5 45.0 76.4 65.5 208.0
20 55.8 45.3 76.8 65.9 189.0
100 55.6 45.2 76.9 65.9 66.4
Figure 10: Mesh estimation results of using different denoising steps $T$ for VMarker-Pro on a sample data from H3.6M [13] dataset. Each hypothesis is denoised from a zero initial noise for a fair comparison.
Denoising steps. Table X shows the metrics on the H3.6M [13] and 3DPW [14] test sets when using different numbers of denoising steps $T$. We report the metrics of predicting $S=10$ hypotheses. We also report the speed of the whole model inferring a hypothesis on an NVIDIA A100-PCIE-80GB GPU. It can be seen that using only $T=1$ denoising step results in significant errors, and the estimated mesh is not human-shaped, as shown in Figure 10. As the number of denoising steps $T$ increases, the accuracy of the estimated mesh gradually improves, and the error decreases. When $T>=10$, the accuracy of the mesh estimation stabilizes, and further steps yield only marginal improvements, as validated by the marginal reduction in error and the indistinguishable differences in visualization. Therefore, based on the acceleration of DDIM [60], we choose $T=10$ steps for a good balance between estimation performance and inference speed.
TABLE XI: Robustness to occlusion for VMarker and VMarker-Pro on 3DPW [14] test set when different body parts are occluded. We report MPVE error.
Occ. Parts VMarker VMarker-Pro
$S=10$ $S=100$
None 77.9 77.3 67.4
2 Arms 79.2 78.7 68.8
2 Legs 78.3 77.6 67.8
Body 78.6 78.1 68.2
Random 78.7 77.9 68.4
Figure 11: Top: Meshes estimated by our deterministic approach VMarker on images from 3DPW test set. The rightmost case in the dashed box shows a typical failure. Bottom: Meshes estimated by VMarker on Internet images with challenging cases (extreme shapes or in a long dress).
Figure 12: (a) For each virtual marker (represented by a star), we highlight the top 30 most affected vertices (represented by a colored dot) based on average coefficient matrix $\mathbf{A}$. (b) For each vertex (dot), we highlight the top 3 virtual markers (star) that contribute the most. The dependency has a strong locality which improves the robustness when some virtual markers cannot be accurately detected.
Figure 13: Qualitative comparison of VMarker and VMarker-Pro on 3DPW [14] testset. VMarker may fail when facing ambiguity, while VMarker-Pro could generate multiple reasonable solutions. For each mesh estimate, we present both a projected view and a side view in each column. We show three hypotheses generated by VMarker-Pro.
Figure 14: Meshes estimated by VMarker-Pro on Internet images. For each case in a row, we present three hypotheses, showing both a projected view and a side view in each column. The rightmost column shows the overlapped side views to unveil the subtle differences.
Figure 15: A typical failure case of VMarker-Pro. The heavy occlusion and ambiguity cause the model to incorrectly estimate the right arm and right leg of the male dancer.
Robustness to occlusion. We report results of VMarker and VMarker-Pro when different image regions of corresponding virtual markers are occluded by a synthetic random mask in Table XI. The errors of VMarker are slightly larger than the original image (None). We further analyze how inaccurate virtual markers would affect the mesh estimation, i.e.  when part of human body is occluded or truncated. According to the finally learned coefficient matrix $\mathbf{A}$ of our model VMarker, we highlight the relationship weights among virtual markers and all vertices in Figure 12. We can see that our model actually learns local and sparse dependency between each vertex and the virtual markers, e.g.  for each vertex, the virtual markers that contribute the most are in a near range as shown in Figure 12 (b). Therefore, in inference, if a virtual marker has inaccurate position estimation due to occlusion or truncation, the dependent vertices may have inaccurate estimates, while the rest will be barely affected.
This locality enhances the robustness of the VMarker-Pro method against occlusion as well. When sampling $S=10$ hypotheses, VMarker-Pro exhibits a minor increase in error under various circumstances of VM occlusion.
4.6 Qualitative Results
VMarker. Figure 11 (top) presents some meshes estimated by our approach VMarker on natural images from the 3DPW test set. The rightmost case shows a typical failure where our method has a wrong pose estimate of the left arm due to heavy occlusion. We can see that the failure is constrained to the local region and the rest of the body still gets accurate estimates. As discussed in Sec. 15, we suppose that this is due to the advantages brought by the locality of the virtual marker representation, which can confine erroneous mesh estimates to a local range and thereby enhance robustness. Figure 2 (right) shows more examples where occlusion or truncation occurs, and our method can still get accurate or reasonable estimates robustly. Note that when truncation occurs, our method still guesses the positions of the truncated virtual markers.
Figure 11 (bottom) shows our estimated meshes on challenging cases, which indicates the strong generalization ability of our model on diverse postures and actions in natural scenes. Please refer to the supplementary for more quality results. Note that since the datasets do not provide supervision of head orientation, face expression, hands, or feet, the estimates of these parts are just in canonical poses inevitably.
VMarker-Pro. Figure 13 compares the visualization of VMarker and VMarker-Pro on the 3DPW [14] test set. We draw three hypotheses generated by VMarker-Pro, depicted in blue. Additionally, we include the results from the deterministic model VMarker (green). It is noticeable that, due to self-occlusion or object occlusion, VMarker might yield results that seem accurate in 2D projections yet are erroneous in 3D space, e.g.  the incorrect estimation of the right leg in the second case. This issue stems from the inherent ambiguity associated with occlusions. By leveraging the advantages of the data distribution learned by the denoising model, VMarker-Pro can generate more reasonable solutions, such as the naturally standing legs.
In Figure 14, we show more results of VMarker-Pro on unseen Internet images. It can be observed that VMarker-Pro possesses strong generalization capabilities, obtaining reasonable 3D estimations when faced with uncertain ambiguities, such as the various plausible poses of the runner’s right arm in the second case (as can be seen from the side view). Furthermore, even under complex actions and backgrounds, such as the person climbing in the third image, VMarker-Pro also demonstrates robust estimation results.
Figure 15 shows a typical failure case of VMarker-Pro on Internet image, where our method fails when heavy occlusion and ambiguity happen. Our method incorrectly estimates the right arm and right leg of the male dancer, which may be addressed using more powerful estimators or more diverse training datasets in the future.
5 Conclusion
In this paper, we present a novel intermediate representation virtual marker, which is more expressive than the prevailing skeleton representation and more accessible than physical markers. Our VMarker method, built on virtual markers, excels in detecting these markers in wild images to interpolate and reconstruct complete, realistically shaped meshes. To further address the ambiguity in 3D virtual marker prediction due to occlusions, we progress to VMarker-Pro, a probabilistic approach that generates multiple plausible meshes in alignment with the image. VMarker-Pro effectively captures the 3D virtual marker distribution through a conditional diffusion process, enabling precise modeling of 3D mesh distributions. Experiments demonstrate the effectiveness of our approaches and they show advanced generalization potential in spite of its simplicity.
Acknowledgments
This work was supported by National Science and Technology Major Project (2022ZD0114904).
Appendix A Post-processing on Virtual Markers
As described in Section 3.1, considering the left-right symmetric human body structure, we slightly adjust the learned virtual markers $\mathbf{Z}$ to be symmetric. In fact, after the first step that updates each $\mathbf{z}_{i}$ by its nearest vertex to get $\widetilde{\mathbf{Z}}\in\mathbb{R}^{3\times K}$. $\widetilde{\mathbf{Z}}$ are almost symmetric with few exceptions. To get the final symmetric virtual markers $\widetilde{\mathbf{Z}}^{sym}\in\mathbb{R}^{3\times K}$, for each virtual marker located in the left body part, we take its symmetric vertex in the right body to be its symmetric counterpart.
Since the human mesh (i.e.  SMPL [12]) itself is not strictly symmetric, we clarify the symmetric vertex pair (e.g.  left elbow and right elbow) on a human mesh template $\mathbf{X}^{t}\in\mathbb{R}^{3\times M}$ in Figure 16. We place $\mathbf{X}^{t}$ at the origin of the 3D coordinate system. Formally, we define the cost of matching $i^{th}$ vertex to $j^{th}$ vertex to be $\bm{C}_{i,j}=\left|x_{i}+x_{j}\right|+\left|y_{i}-y_{j}\right|+\left|z_{i}-z_{j}\right|$. A symmetric vertex pair $(\mathbf{X}^{t}_{i},\mathbf{X}^{t}_{j})$ is defined to have the minimal cost $\bm{C}_{i,j}$. In this way, for each virtual marker in the left body, we take its symmetric vertex counterpart to be its symmetric virtual marker and finally get $\widetilde{\mathbf{Z}}^{sym}$.
Figure 16: Illustration of the human mesh template $\mathbf{X}^{t}$ at the 3D coordinate system and a symmetric vertex pair $(\mathbf{X}^{t}_{i},\mathbf{X}^{t}_{j})$.
Appendix B Experiments
In this section, we first add detailed descriptions for datasets and then provide more experimental results of our approach.
B.1 Datasets
H3.6M [13]. Following previous works [6, 33, 16, 102], we use the SMPL parameters generated from MoSh [5], which are fitted to the 3D physical marker locations, to get the GT 3D mesh supervision. Following standard practice [6], we evaluate the quality of 3D pose of $14$ joints derived from the estimated mesh, i.e.  $\hat{\mathbf{M}}\mathcal{J}$. We report Mean Per Joint Position Error (MPJPE) and PA-MPJPE in millimeters (mm). The latter uses Procrustes algorithm [103] to align the estimates to GT poses before computing MPJPE. To evaluate mesh estimation results, we also report Mean Per Vertex Error (MPVE) which can be interpreted as MPJPE computed over the whole mesh.
3DPW [14]. The 3D GT SMPL parameters are obtained by using the data from IMUs when collected. Following the previous works [87, 89, 47, 40], we use the train set of 3DPW to learn the model and evaluate on the test set. The same evaluation metrics as H3.6M are used. We further evaluate the performance of our probabilistic framework VMarker-Pro on two subsets of 3DPW, i.e.  3DPW-OC and 3DPW-PC [14, 95] which have both object and person occlusion. In this setting, we do not use the 3DPW training set for a fair comparison following [96].
MPI-INF-3DHP [97] is a 3D pose dataset with 3D GT pose annotations. Since this dataset does not provide 3D mesh annotations, following [6, 33], we only enforce supervision on the 3D skeletons (Eq. 9) in mesh losses.
UP-3D [15] is a wild 2D pose dataset with natural images. The 3D poses and meshes are obtained by SMPLify [29]. Due to the lack of GT 3D poses, the fitted meshes are not accurate. Therefore we only use the 2D annotations to train the 3D virtual marker estimation network as in [18].
COCO [98] is a large wild 2D pose dataset with natural images. Previous work [17] used SMPLify-X [30] to obtain pseudo SMPL mesh annotations but they are not accurate. However, we find that if we project the 3D mesh to 2D image, the resulting 2D mesh vertices align well with the image. So we leverage the 2D annotations to train the virtual marker estimation network as in [18].
SURREAL [27] is a large-scale synthetic dataset containing 6 million frames of synthetic humans. The images are photo-realistic renderings of people under large variations in shape, texture, viewpoint, and body pose. To ensure realism, the synthetic bodies are created using the SMPL body model, whose parameters are fit by the MoSh [5] given raw 3D physical marker data. All the images have a resolution of $320\times 240$. We use the same training split to train the model and evaluate the test split following [1].
B.2 Implementation Details
Following common practice [6, 1, 17, 40, 16, 84, 89, 87], we conduct mix-training by using the above 2D and 3D datasets for experiments on the H3.6M and 3DPW datasets. To leverage the 3D pose estimation dataset, i.e.  MPI-INF-3DHP [97], we extend the $64$ virtual markers with the $17$ landmark joints (i.e.  skeleton) from the MPI-INF-3DHP dataset. For experiments on the SURREAL dataset, we use its training set alone as in [1, 36]. We implement the proposed method with PyTorch. The experiments for VMarker and VMarker-Pro are conducted on a Linux machine with 4 NVIDIA 16GB V100 GPUs and 2 NVIDIA A100-PCIE-80G GPUs, respectively.
To train VMarker, the whole network is trained for 40 epochs with batch size 32 using Adam [100] optimizer. We set $\lambda_{3d},\lambda_{c},\lambda_{m},\lambda_{e}=100,0.1,0.1,0.1$ for each loss term. To train VMarker-Pro, the whole network is trained for 50 epochs with batch size 80 using Adam [100] optimizer. We set $\lambda_{diffusion}=1$ and keep the rest loss weights the same as training VMarker. For implementing LCN unit, we modify the original LCN network [77] by injecting a Cross-attention [82] module in each LCN layer. For more implementation details, please check https://github.com/ShirleyMaxx/VMarker-Pro.
B.3 Additional Quantitative Results
Different loss terms. Table XII reports the MPVE error on H3.6M [13] test set when ablating different loss terms. The confidence loss [72] is used to encourage the interpretability of the heatmaps to have a maxima response at the GT position. Without the confidence loss, the error increases slightly. If ablating the surface losses, MPVE increases a lot, which demonstrates the smoothing effect of these two terms.
Comparison to fitting. In order to disentangle the ability of mesh regression from markers using $\hat{\mathbf{A}}$ with the ability to detect the virtual markers accurately from images, we first compute the estimation errors of the virtual markers. The MPJPE over all the virtual markers is $35.5$mm, which demonstrates that these virtual markers can be accurately detected from the images. We then fit the mesh model parameters to these virtual markers. Table XIII shows the metrics of the fitted mesh on the SURREAL [27] test set. As we can see, the fitted mesh has a similar error as our regression ones which uses the interpolation matrix $\hat{\mathbf{A}}$, which validates the accuracy of the estimated virtual markers.
B.4 Additional Qualitative Results
Figure 19 shows more qualitative comparisons with Pose2Mesh [1] on the SURREAL test set in which has diverse body shapes. The skeleton representation used in Pose2Mesh [1] loses the body shape information so the method can only recover mean shapes, regardless of whether the person is slim or stout. This is caused by the limited skeleton representation bottleneck so that the model learns a mean shape for the whole training dataset implicitly. In contrast, both of our approaches VMarker and VMarker-Pro generate more accurate mesh estimation results with the help of virtual marker representation.
TABLE XII: MPVE errors on H3.6M [13] test set when ablating different loss terms.
VMarker w/o w/o w/o w/o
$\mathcal{L}_{conf}$ $\mathcal{L}_{pose}$ $\mathcal{L}_{normal}$ $\mathcal{L}_{edge}$
[…]
TABLE XIII: Results on SURREAL [27] test set when the mesh is obtained by fitting to the estimated virtual markers.
[…]
Fitting 44.6 34.8 29.5
VMarker 44.7 36.9 28.9
Figure 17: Meshes estimated by VMarker on Internet images with challenging cases (complex poses or extreme body shapes).
Figure 18: Meshes estimated by VMarker-Pro on Internet images with challenging cases (complex poses or extreme body shapes). For each case in a row, we present three hypotheses, showing both a projected view and a side view in each column. The rightmost column shows the overlapped side views to unveil the subtle differences.
Figure 20 shows more qualitative comparisons with Pose2Mesh [1] and METRO [87] on the 3DPW test set. Pose2Mesh and METRO use the skeleton or all 3D vertices as intermediate representations, respectively. The estimated meshes are overlaid on the images according to the camera parameters. Pose2Mesh [1] has difficulty in estimating correct body pose and shapes when truncation occurs (a) or in complex postures (c). The results of METRO [87] have many artifacts where the estimated mesh is not smooth, and they also fail to align the image well. In contrast, our method estimates more accurate human poses and shapes and has smooth human mesh results. In addition, it is more robust to truncation and occlusion and aligns the image better.
Figure 17 and 18 show more qualitative results of VMarker and VMarker-Pro on Internet images with challenging cases, respectively. These images include humans with diverse body shapes or complex poses. Our methods generalize well on the natural scenes. When occlusion happens, the probabilistic method VMarker-Pro could infer multiple reasonable poses, e.g.  the different angles of the left arm of the running woman in the first case in Figure 18.
Figure 19: Qualitative comparison between our methods and Pose2Mesh [1] on SURREAL test set [27]. Our approaches generate more accurate mesh estimation results on images of diverse body shapes. We show 2 hypotheses generated by VMarker-Pro.
Figure 20: Qualitative comparison between our method VMarker and Pose2Mesh [1], METRO [87] on 3DPW test set [14]. Our approach is more robust to occlusion and truncation and generates more accurate mesh estimation results that align images well.
Appendix C Human Subject Data
We use existing public datasets of human subjects in our experiments following their official licensing requirements. With proper usage, the proposed method could be beneficial to society (e.g.  elderly fall detection).

### v4

@@
-VMarker-Pro: Probabilistic 3D Human Mesh Estimation from Virtual Markers
+3D Human Mesh Estimation from Virtual Markers
 Abstract
-Monocular 3D human mesh estimation faces challenges due to depth ambiguity and the complexity of mapping images to complex parameter spaces.
-Recent methods propose to use 3D poses as a proxy representation, which often lose crucial body shape information, leading to mediocre performance.
-Conversely, advanced motion capture systems, though accurate, are impractical for markerless wild images.
-Addressing these limitations, we introduce an innovative intermediate representation as virtual markers, which are learned from large-scale mocap data, mimicking the effects of physical markers.
-Building upon virtual markers, we propose VMarker, which detects virtual markers from wild images, and the intact mesh with realistic shapes can be obtained by simply interpolation from these markers.
-To address occlusions that obscure 3D virtual marker estimation, we further enhance our method with VMarker-Pro, a probabilistic framework that generates multiple plausible meshes aligned with images.
-This framework models the 3D virtual marker estimation as a conditional denoising process, enabling robust 3D mesh estimation.
-Our approaches surpass existing methods on three benchmark datasets, particularly demonstrating significant improvements on the SURREAL dataset, which features diverse body shapes.
-Additionally, VMarker-Pro excels in accurately modeling data distributions, significantly enhancing performance in occluded scenarios.
-Code and models are available at https://github.com/ShirleyMaxx/VMarker-Pro.
-Index Terms:
-Computer vision, 3D human mesh estimation, probabilistic modeling
-Figure 1: Top: Mesh estimation results on examples with different body shapes.
+Inspired by the success of volumetric 3D pose estimation, some recent human mesh estimators propose to estimate 3D skeletons as intermediate representations, from which, the dense 3D meshes are regressed by exploiting the mesh topology.
+However, body shape information is lost in extracting skeletons, leading to mediocre performance.
+The advanced motion capture systems solve the problem by placing dense physical markers on the body surface, which allows to extract realistic meshes from their non-rigid motions.
+However, they cannot be applied to wild images without markers.
+In this work, we present an intermediate representation, named virtual markers, which learns 64 landmark keypoints on the body surface based on the large-scale mocap data in a generative style, mimicking the effects of physical markers.
+The virtual markers can be accurately detected from wild images and can reconstruct the intact meshes with realistic shapes by simple interpolation.
+Our approach outperforms the state-of-the-art methods on three datasets.
+In particular, it surpasses the existing methods by a notable margin on the SURREAL dataset, which has diverse body shapes.
+Code is available at https://github.com/ShirleyMaxx/VirtualMarker.
+1 Introduction
+Figure 1: Mesh estimation results on four examples with different body shapes.
 Pose2Mesh [7] which uses 3D skeletons as the intermediate representation fails to predict accurate shapes.
-Our virtual marker-based method VMarker obtains accurate estimates.
-Bottom: VMarker may fail when occlusion happens, e.g.
-the occluded left arm is wrongly estimated.
-To handle the ambiguity, we propose a probabilistic framework VMarker-Pro which could generate multiple reasonable estimates aligning with the image.
-1 Introduction
-Monocular 3D human mesh estimation aims to estimate 3D positions of the mesh vertices that are on the body surface from an RGB image.
-The task has attracted a lot of attention from the computer vision and computer graphics communities [1, 2, 3, 4, 5, 6, 7, 8] because it can benefit many applications such as virtual reality [9].
-Recently, deep learning-based methods [6, 1, 10] have significantly advanced the accuracy on benchmark datasets.
-Pioneer methods [11, 6] propose to regress the pose and shape parameters of the mesh models such as SMPL [12] directly from images, learning a one-to-one mapping from images to the parameter space.
-While straightforward, their performance usually suffers from several limitations.
-Firstly, the mapping from the image to the $SO(3)$ parameter space is highly non-linear and suffers from image-model misalignment [10].
-Secondly, existing mesh datasets [13, 14, 15] are small and limited to simple laboratory environments due to the complex capturing process.
+Our virtual marker-based method obtains accurate estimates.
+3D human mesh estimation aims to estimate the 3D positions of the mesh vertices that are on the body surface.
+The task has attracted a lot of attention from the computer vision and computer graphics communities [3, 43, 30, 35, 51, 19, 25, 37, 27, 10] because it can benefit many applications such as virtual reality [15].
+Recently, the deep learning-based methods [19, 7, 29] have significantly advanced the accuracy on the benchmark datasets.
+The pioneer methods [51, 19] propose to regress the pose and shape parameters of the mesh models such as SMPL [36] directly from images.
+While straightforward, their accuracy is usually lower than the state-of-the-arts.
+The first reason is that the mapping from the image features to the model parameters is highly non-linear and suffers from image-model misalignment [29].
+Besides, existing mesh datasets [16, 54, 38, 28] are small and limited to simple laboratory environments due to the complex capturing process.
 The lack of sufficient training data severely limits its performance.
-Lastly, due to the inherent depth ambiguity in monocular estimation, multiple 3D meshes may correspond to the same 2D projection.
-As a result, these deterministic methods may produce incorrect estimates in the presence of occlusions.
 Recently, some works [26, 39] begin to formulate mesh estimation as a dense 3D keypoint detection task inspired by the success of volumetric pose estimation [47, 50, 65, 44, 59, 45].
@@
 However, it is computationally expensive because it has more than several thousand vertices.
-Moon et al.
-[17] improves the efficiency by decomposing the 3D heatmaps into multiple 1D heatmaps at the cost of mediocre accuracy.
+Moon and Lee [39] improve the efficiency by decomposing the 3D heatmaps into multiple 1D heatmaps at the cost of mediocre accuracy.
 Choi et al.
-[1] propose to first detect a sparser set of skeleton joints in the images, from which dense 3D meshes are regressed by exploiting the mesh topology.
+[7] propose to first detect a sparser set of skeleton joints in the images, from which the dense 3D meshes are regressed by exploiting the mesh topology.
 The methods along this direction have attracted increasing attention [7, 29, 55] due to two reasons.
@@
 Second, mesh regression from the skeletons is efficient.
-However, important information about the body shapes is lost in extracting the 3D skeletons, which is largely overlooked.
-As a result, different types of body shapes, whether lean or obese, cannot be accurately estimated, e.g.
-the results of Pose2Mesh [1] in Figure 1.
-The professional marker-based motion capture method MoSh [5] places physical markers on the body surface to extract meshes with accurate shapes under non-rigid motions.
-However, the physical markers limit the approach in laboratory environments.
-We are inspired to think whether we can identify a set of landmarks on the mesh as virtual markers, e.g.
-, elbow and wrist, that can be detected from wild images, allowing us to recover accurate body shapes.
+However, important information about the body shapes is lost in extracting the 3D skeletons, which is largely overlooked previously.
+As a result, different types of body shapes, such as lean or obese, cannot be accurately estimated (see Figure 1).
+The professional marker-based motion capture (mocap) method MoSh [35] places physical markers on the body surface and explore their subtle non-rigid motions to extract meshes with accurate shapes.
+However, the physical markers limit the approach to be used in laboratory environments.
+We are inspired to think whether we can identify a set of landmarks on the mesh as virtual markers, e.g., elbow and wrist, that can be detected from wild images, and allow to recover accurate body shapes?
 The desired virtual markers should satisfy several requirements.
-First, the number of markers should be much smaller than that of the mesh vertices so that we can use volumetric representations to estimate their 3D positions efficiently.
+First, the number of markers should be much smaller than that of the mesh vertices so that we can use volumetric representations to efficiently estimate their 3D positions.
 Second, the markers should capture the mesh topology so that the intact mesh can be accurately regressed from them.
-Third, the virtual markers have distinguishable visual patterns and can be detected from images.
+Third, the virtual markers have distinguishable visual patterns so that they can be detected from images.
 In this work, we present a learning algorithm based on archetypal analysis [12] to identify a subset of mesh vertices as the virtual markers that try to satisfy the above requirements to the best extent.
 Figure 2 shows that the learned virtual markers coarsely outline the body shape and pose which paves the way for estimating meshes with accurate shapes.
-Then we present a simple framework VMarker for 3D mesh estimation on top of the representation as shown in Figure 3 (top).
-It first detects the 3D positions of the virtual markers [18].
+Then we present a simple framework for 3D mesh estimation on top of the representation as shown in Figure 3.
+It first learns a 3D keypoint estimation network based on [47] to detect the 3D positions of the virtual markers.
 Then we recover the intact mesh simply by interpolating them.
 The interpolation weights are pre-trained in the representation learning step and will be adjusted by a light network based on the prediction confidences of the virtual markers for each image.
-Although the virtual marker representation eases the mesh estimation task, the vulnerability of deterministic methods to depth ambiguity and occlusions underscores the need for probabilistic modeling.
-For example, in Figure 1 (bottom), VMarker wrongly estimates the occluded left arm.
-Expanding upon VMarker, we further develop VMarker-Pro, a probabilistic framework capable of estimating multiple plausible 3D meshes that are consistent with the given 2D observation, as depicted in Figure 3 (bottom).
-Motivated by recent breakthroughs in generative models, we utilize diffusion models [26] to enhance the diversity and robustness of our mesh predictions.
-Concretely, we formulate the estimation of 3D virtual markers as a conditional reverse denoising process.
-Based on the input 2D image and the predicted virtual markers from VMarker, we probabilistically model the 3D marker positions and finally obtain multiple reasonable 3D mesh estimates.
-For instance, VMarker-Pro could estimate a more naturally swinging left arm during walking than VMarker as shown in the bottom of Figure 1.
-We extensively evaluate our approaches on three benchmark datasets.
-Both VMarker and VMarker-Pro consistently outperforms the state-of-the-art methods across the benchmarks.
-Notably, VMarker demonstrates remarkable gains on the SURREAL dataset [27], which has a variety of body shapes.
-Ablation studies further affirm the merits of virtual marker representation, particularly in estimating accurate shapes.
-Furthermore, VMarker-Pro stands out in its precise modeling of data distributions, markedly enhancing performance in occlusion-heavy scenarios.
-Finally, experiments show decent generalization ability of our methods which generate visually appealing results for wild images.
-A preliminary version of this work was presented in CVPR 2023 [28].
-Our new contributions are three-fold.
-First, we propose VMarker-Pro, which incorporates probabilistic modeling to handle the inherent ambiguity in monocular 3D mesh estimation.
-This yields a more flexible and robust human mesh distribution estimation.
-Second, VMarker-Pro outperforms state-of-the-arts (SOTA) on multiple benchmark datasets in terms of the accuracy of distribution.
-Finally, VMarker-Pro significantly improves the performance in scenarios with occlusions than VMarker and other SOTA methods, demonstrating its ability to resolve ambiguity.
+We extensively evaluate our approach on three benchmark datasets.
+It consistently outperforms the state-of-the-art methods on all of them.
+In particular, it achieves a significant gain on the SURREAL dataset [53] which has a variety of body shapes.
+Our ablation study also validates the advantages of the virtual marker representation in terms of recovering accurate shapes.
+Finally, the method shows decent generalization ability and generates visually appealing results for the wild images.
 2 Related work
@@
 For example, Loper et al.
-[5] propose MoSh that optimizes the SMPL [12] parameters to align the mesh with the 3D marker positions.
+[35] propose MoSh that optimizes the SMPL parameters to align the mesh with the 3D marker positions.
 It is usually used to get GT 3D meshes for benchmark datasets because of its high accuracy.
@@
 The virtual markers act similarly to physical body markers and approximately outline the body shape.
-Right: Mesh estimation results by our approach VMarker, from left to right are input image, estimated 3D mesh overlayed on the image, and three different viewpoints showing the estimated 3D mesh with our intermediate predicted virtual markers (blue balls), respectively.
+Right: Mesh estimation results by our approach, from left to right are input image, estimated 3D mesh overlayed on the image, and three different viewpoints showing the estimated 3D mesh with our intermediate predicted virtual markers (blue balls), respectively.
 2.2 Learning-based mesh estimation
-Recently, most studies follow the learning-based framework and achieve promising results.
-Existing learning-based methods can roughly be divided into two categories: deterministic mesh estimation and probabilistic mesh estimation.
-A considerable amount of deterministic estimation work [11, 6, 33, 34, 35, 36, 17, 37, 38, 39, 40, 10, 1] aims at estimating a single 3D human mesh from a monocular image.
-However, considering the inherent depth ambiguity in monocular settings, an increasing number of works propose the probabilistic modeling framework which estimates multiple feasible solutions [41, 42, 43, 44, 45] to enhance the robustness of the model.
-2.2.1 Deterministic mesh estimation
-Pioneer works propose to use deep networks [11, 6, 33, 36] to regress the SMPL [12] parameters from image features.
+Recently, most works follow the learning-based framework and have achieved promising results.
+Deep networks [51, 19, 25, 37, 27] are used to regress the SMPL parameters from image features.
 However, learning the mapping from the image space to the parameter space is highly non-linear [39].
@@
 However, accurate body shapes are difficult to be recovered from the oversimplified 3D skeletons.
-2.2.2 Probabilistic mesh estimation
-Despite the extensive exploration of multi-hypothesis 3D pose estimation [49, 50, 51, 52, 53, 54, 55], research on probabilistic 3D human mesh estimation remains relatively underdeveloped.
-Most recent approaches [41, 56, 42, 44, 45] predict distributions over SMPL [12] parameters conditioned on a 2D image.
-Biggs et al.
-[41] expand upon HMR [6] to predict a categorical distribution over SMPL parameters and obtain multiple hypotheses discreetly.
-Sengupta et al.
-[56] and ProPose [42] utilize the matrix Fisher distribution to model body pose rotation distributions.
-ProHMR [44] and HuManiFlow [43] employ Normalizing Flows [57, 58] to model the feasible 3D human model parameter distribution.
-More recently, HMDiff [45] adopts diffusion models [26] to generate multiple plausible human meshes.
-Our work belongs to the learning-based class, offering both deterministic and probabilistic approaches.
-The deterministic approach, VMarker, introduces virtual markers as a novel intermediate representation, improving upon traditional physical markers or skeletons to better disambiguate pose and shape with more expressiveness.
-Our probabilistic framework, VMarker-Pro, employs diffusion models to produce diverse and plausible virtual marker estimates, from which multiple full mesh estimates that align with the image cues can be obtained.
-2.3 Diffusion models
-Denoising Diffusion Probabilistic Models (DDPMs) [59, 26] are a class of generative models that garner attention for their ability to generate high-quality images [26].
-DDPMs employ a forward process where noise is incrementally added to the data until a noise distribution is reached, paired with a reverse process that incrementally denoises to generate data from noise.
-The strength of DDPMs lies in their iterative refinement, effectively capturing the target data distribution and enabling the synthesis of new samples.
-With the development of acceleration [26, 60] and enhancement [61, 62], the versatility of DDPMs has been demonstrated across various tasks beyond image generation [26, 60], contributing to innovations in video generation [63, 64, 65], semantic segmentation [66], motion generation [67, 68, 69], etc.
-In the realm of 3D mesh estimation, existing methods [45] apply DDPMs to model a set of downsampled mesh vertices, which is hard to detect from images [17, 28].
-Our approach utilizes DDPMs to model the distribution of virtual markers, an efficient and expressive intermediate representation.
-This reduces the complexity of the modeling, resulting in more robust and accurate estimations.
+Our work belongs to the learning-based class and is related to works that use physical markers or skeletons as intermediate representations.
+But different from them, we propose a novel intermediate representation, named virtual markers, which is more expressive to reduce the ambiguity in pose and shape estimation than body skeletons and can be applied to wild images.
 3 Method
@@
 First, Section 3.1 introduces how we learn the virtual marker representation from mocap data.
-Based on the virtual marker, we present the deterministic framework VMarker for mesh estimation from an image and its training details in Section 3.2.
-Then we extend VMarker to a probabilistic mesh estimation framework, VMarker-Pro in Section 3.3 and introduce its training details as well as the inference strategy.
+Then we present the overall framework for mesh estimation from an image in Section 3.2.
+At last, Section 3.3 discusses the loss functions and training details.
 3.1 The virtual marker representation
-We represent a 3D mesh by a vector of vertex positions $\mathbf{x}\in\mathbb{R}^{3M}$ where $M$ is the number of mesh vertices.
+We represent a mesh by a vector of vertex positions $\mathbf{x}\in\mathbb{R}^{3M}$ where $M$ is the number of mesh vertices.
 Denote a mocap dataset such as [16] with $N$ meshes as $\overset{\frown}{\mathbf{X}}=[\mathbf{x}_{1},\,...,\,\mathbf{x}_{N}]\in\mathbb{R}^{3M\times N}$.
 To unveil the latent structure among vertices, we reshape it to $\mathbf{X}\in\mathbb{R}^{3N\times M}$ with each column $\mathbf{x}_{i}\in\mathbb{R}^{3N}$ representing all possible positions of the $i^{\text{th}}$ vertex in the dataset [16].
-The mesh $\mathbf{X}$ can be approximated by a low-rank matrix because the mesh representation is smooth and redundant where some vertices can be accurately reconstructed by others.
-Inspired by this, we aim to learn $K$ virtual markers from $M$ mesh vertices $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{K}]\in\mathbb{R}^{3N\times K}$, as shown in Figure 2, that try to satisfy the following two requirements to the greatest extent.
+The rank of $\mathbf{X}$ is smaller than $M$ because the mesh representation is smooth and redundant where some vertices can be accurately reconstructed by the others.
+While it seems natural to apply PCA [18] to $\mathbf{X}$ to compute the eigenvectors as virtual markers for reconstructing others, there is no guarantee that the virtual markers correspond to the mesh vertices, making them difficult to be detected from images.
+Instead, we aim to learn $K$ virtual markers $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{K}]\in\mathbb{R}^{3N\times K}$ that try to satisfy the following two requirements to the greatest extent.
 First, they can accurately reconstruct the intact mesh $\mathbf{X}$ by their linear combinations: $\mathbf{X}=\mathbf{Z}\mathbf{A}$, where $\mathbf{A}\in\mathbb{R}^{K\times M}$ is a coefficient matrix that encodes the spatial relationship between the virtual markers and the mesh vertices.
@@
 Ideally, they can be on the body surface as the meshes.
-While it seems natural to apply singular value decomposition (SVD) [70] to $\mathbf{X}$ to compute the eigenvectors as virtual markers for reconstructing others, there is no guarantee that the virtual markers correspond to the mesh vertices, making them difficult to be detected from images.
-To that end, we apply archetypal analysis [25, 71] to learn $\mathbf{Z}$ by minimizing a reconstruction error with two additional constraints: (1) each vertex $\mathbf{x}_{i}$ can be reconstructed by convex combinations of $\mathbf{Z}$, and (2) each marker $\mathbf{z}_{i}$ should be convex combinations of the mesh vertices $\mathbf{X}$:
+We apply archetypal analysis [12, 4] to learn $\mathbf{Z}$ by minimizing a reconstruction error with two additional constraints: (1) each vertex $\mathbf{x}_{i}$ can be reconstructed by convex combinations of $\mathbf{Z}$, and (2) each marker $\mathbf{z}_{i}$ should be convex combinations of the mesh vertices $\mathbf{X}$:
 […]
 where $\mathbf{A}=[\bm{\alpha}_{1},...,\bm{\alpha}_{M}]\in\mathbb{R}^{K\times M}$, each $\bm{\alpha}$ resides in the simplex ${\Delta}_{K}\triangleq\{\bm{\alpha}\in\mathbb{R}^{K}\,\mathrm{s.t.}\,\bm{\alpha}\succeq 0\,\text{and}\,{||\bm{\alpha}||}_{1}=1\}$, and $\mathbf{B}=[\bm{\beta}_{1},...,\bm{\beta}_{K}]\in\mathbb{R}^{M\times K}$, $\bm{\beta}_{j}\in{\Delta}_{M}$.
-We adopt Active-set algorithm [71] to solve objective (Eq.
-1) and obtain the learned virtual markers $\mathbf{Z}=\mathbf{X}\mathbf{B}\in\mathbb{R}^{3N\times K}$.
+We adopt Active-set algorithm [4] to solve objective (1) and obtain the learned virtual markers $\mathbf{Z}=\mathbf{X}\mathbf{B}\in\mathbb{R}^{3N\times K}$.
 As shown in [12, 4], the two constraints encourage the virtual markers $\mathbf{Z}$ to unveil the latent structure among vertices, therefore they learn to be close to the extreme points of the mesh and located on the body surface as much as possible.
-Figure 3: Overview of our deterministic (top) and probabilistic (bottom) framework.
-Top: The deterministic framework VMarker.
-Given an input image, it first estimates the 3D virtual marker (VM) positions $\mathbf{P}$.
-Then we update the coefficient matrix $\mathbf{A}$ based on the estimation confidence scores $\mathbf{C}$ of the VM.
-Finally, the complete human mesh can be simply recovered by linear multiplication $\mathbf{M}=\mathbf{P}\mathbf{A}$.
-Bottom: The probabilistic framework VMarker-Pro.
-Building upon VMarker  we formulate the 3D VM estimation as a reverse denoising process using diffusion models.
-We introduce a denoiser that learns to remove the added noise from a noisy 3D VM $\hat{\mathbf{P}}(t)$ during training, conditioned on the 2D image feature $\mathbf{F}$ and the estimated 2D VMs $\mathbf{P}^{2d}$, i.e.
-the first two dimensions of $\mathbf{P}$.
-At inference, we iteratively removes noise from random noise $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ according to Eq.
-16, yielding a denoised 3D VMs $\mathbf{P}(0)$.
-Finally, we obtain the estimated mesh $\mathbf{M}=\mathbf{P}(0)\mathbf{A}$.
-VMarker-Pro can generate multiple feasible solutions aligned with 2D image cues during inference.
+Figure 3: Overview of our framework.
+Given an input image $\mathbf{I}$, it first estimates the 3D positions $\hat{\mathbf{P}}$ of the virtual markers.
+Then we update the coefficient matrix $\hat{\mathbf{A}}$ based on the estimation confidence scores $\mathbf{C}$ of the virtual markers.
+Finally, the complete human mesh can be simply recovered by linear multiplication $\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}$.
+Type Formula Reconst. Error (mm) $\downarrow$
+Original $||\mathbf{X}-\mathbf{X}\mathbf{B}\mathbf{A}||^{2}_{F}$ 11.67
+Symmetric $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ 10.98
+Table 1: The reconstruction errors using the original and the symmetric sets of markers on the H3.6M dataset [16], respectively.
+The errors are small indicating that they are sufficiently expressive and can reconstruct all vertices accurately.
 Post-processing.
@@
 We first replace each $\mathbf{z}_{i}\in\mathbf{Z}$ by its nearest vertex on the mesh and obtain $\widetilde{\mathbf{Z}}\in\mathbb{R}^{3\times K}$.
-This is done on a template mesh which is in a canonical pose.
 This step allows us to compute the left or right counterpart of each marker.
 Then we replace the markers in the right body with the symmetric vertices in the left body and obtain the symmetric markers $\widetilde{\mathbf{Z}}^{sym}\in\mathbb{R}^{3\times K}$.
-Finally, we update $\mathbf{B}$ and $\mathbf{A}$ by minimizing $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ subject to $\widetilde{\mathbf{Z}}^{sym}=\mathbf{X}\widetilde{\mathbf{B}}^{sym}$ over all $N$ mesh data.
+Finally we update $\mathbf{B}$ and $\mathbf{A}$ by minimizing $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ subject to $\widetilde{\mathbf{Z}}^{sym}=\mathbf{X}\widetilde{\mathbf{B}}^{sym}$.
 More details are elaborated in the supplementary.
@@
 They are roughly evenly distributed on the surface of the body, and some of them are located close to the body keypoints, which have distinguishable visual patterns to be accurately detected.
-Table I shows the reconstruction errors of using original markers $\mathbf{X}\mathbf{B}$ and the symmetric markers $\mathbf{X}\widetilde{\mathbf{B}}^{sym}$.
+Table 1 shows the reconstruction errors of using original markers $\mathbf{X}\mathbf{B}$ and the symmetric markers $\mathbf{X}\widetilde{\mathbf{B}}^{sym}$.
 Both can reconstruct meshes accurately.
-TABLE I: The reconstruction errors using the original and the symmetric sets of markers on the H3.6M dataset [13], respectively.
-The errors are small indicating that they are sufficiently expressive and can reconstruct all vertices accurately.
-Type Formula Reconst. Error (mm) $\downarrow$
-Original $||\mathbf{X}-\mathbf{X}\mathbf{B}\mathbf{A}||^{2}_{F}$ 11.67
-Symmetric $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ 10.98
-3.2 Deterministic mesh estimation framework
-3.2.1 VMarker
-On top of the virtual markers, we present a simple yet effective framework VMarker for end-to-end 3D human mesh estimation from a single image.
-As shown in Figure 3 (top), it consists of two branches (yellow and orange blocks).
-The first branch (yellow block) uses a volumetric CNN [18] to estimate the 3D virtual marker (VM) positions $\mathbf{P}$, and the second branch (orange block) reconstructs the full mesh $\mathbf{M}$ by predicting a coefficient matrix $\mathbf{A}$:
-$\mathbf{M}=\mathbf{P}\mathbf{A}.$ (2)
+3.2 Mesh estimation framework
+On top of the virtual markers, we present a simple yet effective framework for end-to-end 3D human mesh estimation from a single image.
+As shown in Figure 3, it consists of two branches.
+The first branch uses a volumetric CNN [47] to estimate the 3D positions $\hat{\mathbf{P}}$ of the markers, and the second branch reconstructs the full mesh $\hat{\mathbf{M}}$ by predicting a coefficient matrix $\hat{\mathbf{A}}$:
+$\small\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}.$ (2)
 We will describe the two branches in more detail.
-3D VM estimation.
-We train a neural network to estimate a 3D heatmap $\mathbf{H}=[\mathbf{H}_{1},\,...,\,\mathbf{H}_{K}]\in\mathbb{R}^{K\times D\times H\times W}$ from an image.
+3D marker estimation.
+We train a neural network to estimate a 3D heatmap $\hat{\mathbf{H}}=[\hat{\mathbf{H}}_{1},\,...,\,\hat{\mathbf{H}}_{K}]\in\mathbb{R}^{K\times D\times H\times W}$ from an image.
 The heatmap encodes per-voxel likelihood of each marker.
 There are $D\times H\times W$ voxels in total which are used to discretize the 3D space.
-The 3D position $\mathbf{P}_{z}\in\mathbb{R}^{3}$ of each marker is computed as the center of mass of the corresponding heatmap $\mathbf{H}_{z}$ [18] as follows:
-$\mathbf{P}_{z}=\sum_{d=1}^{D}\sum_{h=1}^{H}\sum_{w=1}^{W}(d,h,w)\cdot\mathbf{H}_{z}(d,h,w).$ (3)
-The positions of all markers are represented as $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\cdots,\mathbf{P}_{K}]$.
-3D mesh estimation.
-Ideally, if we have accurate estimates for all virtual markers $\mathbf{P}$, then we can recover the complete mesh by simply multiplying $\mathbf{P}$ with a fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ with sufficient accuracy as validated in Table I.
+The 3D position $\hat{\mathbf{P}}_{z}\in\mathbb{R}^{3}$ of each marker is computed as the center of mass of the corresponding heatmap $\hat{\mathbf{H}}_{z}$ [47] as follows:
+$\small\hat{\mathbf{P}}_{z}=\sum_{d=1}^{D}\sum_{h=1}^{H}\sum_{w=1}^{W}(d,h,w)\cdot\hat{\mathbf{H}}_{z}(d,h,w).$ (3)
+The positions of all markers are represented as $\hat{\mathbf{P}}=[\hat{\mathbf{P}}_{1},\hat{\mathbf{P}}_{2},\cdots,\hat{\mathbf{P}}_{K}]$.
+Interpolation.
+Ideally, if we have accurate estimates for all virtual markers $\hat{\mathbf{P}}$, then we can recover the complete mesh by simply multiplying $\hat{\mathbf{P}}$ with a fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ with sufficient accuracy as validated in Table 1.
 However, in practice, some markers may have large estimation errors because they may be occluded in the monocular setting.
@@
 Our solution is to rely more on those accurately detected markers.
-To that end, we propose to update the coefficient matrix based on the estimation confidence scores $\mathbf{C}$ of the markers.
-In practice, we simply take the heatmap score at the estimated positions of each marker obtained by interpolation.
-For convenience, we denote this process as $\mathbf{C}_{z}=\mathbf{H}_{z}(\mathbf{P}_{z})$.
-We feed the confidence scores to an updating network which is a single fully-connected (FC) layer to obtain the coefficient matrix $\mathbf{A}$.
-Then the mesh is reconstructed by linear multiplication $\mathbf{M}=\mathbf{P}\mathbf{A}$.
-3.2.2 Training
+To that end, we propose to update the coefficient matrix based on the estimation confidence scores of the markers.
+In practice, we simply take the heatmap score at the estimated positions of each marker, i.e.
+$\hat{\mathbf{H}}_{z}(\hat{\mathbf{P}}_{z})$, and feed them to a single fully-connected layer to obtain the coefficient matrix $\hat{\mathbf{A}}$.
+Then the mesh is reconstructed by $\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}$.
+3.3 Training
 We train the whole network end-to-end in a supervised way. The overall loss function is defined as:
-$\mathcal{L}_{vm}=\lambda_{3d}\mathcal{L}_{3d}+\lambda_{c}\mathcal{L}_{conf}+\lambda_{m}\mathcal{L}_{mesh}.$ (4)
-Virtual marker loss. We define $\mathcal{L}_{3d}$ as the $L_{1}$ distance between the predicted 3D virtual markers $\mathbf{P}$ and the GT $\hat{\mathbf{P}}$ as follows:
-$\mathcal{L}_{3d}=\|\mathbf{P}-\hat{\mathbf{P}}\|_{1}.$ (5)
-Note that it is easy to get GT markers $\hat{\mathbf{P}}$ from GT meshes as stated in Section 3.1 without additional manual annotations.
+$\mathcal{L}=\lambda_{vm}\mathcal{L}_{vm}+\lambda_{c}\mathcal{L}_{conf}+\lambda_{m}\mathcal{L}_{mesh}.$ (4)
+Virtual marker loss. We define $\mathcal{L}_{vm}$ as the $L_{1}$ distance between the predicted 3D virtual markers $\hat{\mathbf{P}}$ and the GT $\hat{\mathbf{P}}^{*}$ as follows:
+$\mathcal{L}_{vm}=\|\hat{\mathbf{P}}-\hat{\mathbf{P}}^{*}\|_{1}.$ (5)
+Note that it is easy to get GT markers $\hat{\mathbf{P}}^{*}$ from GT meshes as stated in Section 3.1 without additional manual annotations.
 Confidence loss.
-We also require that the 3D heatmaps have reasonable shapes, therefore, the heatmap score at the voxel containing the GT marker position $\hat{\mathbf{P}}_{z}$ should have the maximum value as in the previous work [72]:
-$\mathcal{L}_{conf}=-\sum_{z=1}^{K}log(\mathbf{H}_{z}(\hat{\mathbf{P}}_{z})).$ (6)
+We also require that the 3D heatmaps have reasonable shapes, therefore, the heatmap score at the voxel containing the GT marker position $\hat{\mathbf{P}}_{z}^{*}$ should have the maximum value as in the previous work [17]:
+$\mathcal{L}_{conf}=-\sum_{z=1}^{K}log(\hat{\mathbf{H}}_{z}(\hat{\mathbf{P}}_{z}^{*})).$ (6)
 Mesh loss. Following [39], we define $\mathcal{L}_{mesh}$ as a weighted sum of four losses:
-$\mathcal{L}_{mesh}=\mathcal{L}_{vertex}+\mathcal{L}_{pose}+\mathcal{L}_{normal}+\lambda_{e}\mathcal{L}_{edge}.$ (7)
-  • Vertex coordinate loss. We adopt $L_{1}$ loss between predicted 3D mesh coordinates $\mathbf{M}$ with GT mesh $\hat{\mathbf{M}}$ as:
-    $\mathcal{L}_{vertex}=\|\mathbf{M}-\hat{\mathbf{M}}\|_{1}.$ (8)
-  • Pose loss. We use $L_{1}$ loss between the 3D landmark joints regressed from mesh $\mathbf{M}\mathcal{J}$ and the GT joints $\hat{\mathbf{M}}\mathcal{J}$ as:
-    $\mathcal{L}_{pose}=\|\mathbf{M}\mathcal{J}-\hat{\mathbf{M}}\mathcal{J}\|_{1},$ (9)
-    where $\mathcal{J}\in\mathbb{R}^{M\times J}$ is a pre-defined joint regression matrix in SMPL model [29] and $J$ is the number of joints.
+$\small\mathcal{L}_{mesh}=\mathcal{L}_{vertex}+\mathcal{L}_{pose}+\mathcal{L}_{normal}+\lambda_{e}\mathcal{L}_{edge}.$ (7)
+  • Vertex coordinate loss. We adopt $L_{1}$ loss between predicted 3D mesh coordinates $\hat{\mathbf{M}}$ with GT mesh $\hat{\mathbf{M}}^{*}$ as:
+    $\mathcal{L}_{vertex}=\|\hat{\mathbf{M}}-\hat{\mathbf{M}}^{*}\|_{1}.$ (8)
+  • Pose loss. We use $L_{1}$ loss between the 3D landmark joints regressed from mesh $\hat{\mathbf{M}}\mathcal{J}$ and the GT joints $\hat{\mathbf{J}}^{*}$ as:
+    $\mathcal{L}_{pose}=\|\hat{\mathbf{M}}\mathcal{J}-\hat{\mathbf{J}}^{*}\|_{1},$ (9)
+    where $\mathcal{J}\in\mathbb{R}^{M\times J}$ is a pre-defined joint regression matrix in SMPL model [2].
   • Surface losses.
 To improve surface smoothness [56], we supervise the normal vector of a triangle face with GT normal vectors by $\mathcal{L}_{normal}$ and the edge length of the predicted mesh with GT length by $\mathcal{L}_{edge}$:
-    $\mathcal{L}_{normal}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\left<\frac{\mathbf{M}_{i}-\mathbf{M}_{j}}{\|\mathbf{M}_{i}-\mathbf{M}_{j}\|_{2}},\hat{\mathbf{n}}_{f}\right>\right|,$ (10)
-    $\mathcal{L}_{edge}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\|\mathbf{M}_{i}-\mathbf{M}_{j}\|_{2}-\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}\right|.$
-    where $f$ and $\hat{\mathbf{n}}_{f}$ denote a triangle face in the mesh and its GT unit normal vector, respectively.
-$\mathbf{M}_{i}$ denote the $i^{th}$ vertex of $\mathbf{M}$.
-$\hat{}$ denotes GT.
-3.3 Probabilistic mesh estimation framework
-3.3.1 VMarker-Pro
-As shown in Figure 2, virtual markers are often self-occluded, thus there may exist multiple viable solutions.
-An incorrect estimation may introduce errors into the final mesh estimate.
-Therefore, we propose a probabilistic framework for modeling the estimation of 3D VMs, as illustrated in Figure 3 (bottom).
-By leveraging the powerful modeling capabilities of diffusion models [26, 74] on data distributions, we generate multiple feasible 3D VM estimates given an image condition.
-Specifically, we model the estimation of 3D VMs as a conditional reverse denoising process that starts with Gaussian noise and progressively denoises it to obtain the 3D VM using a denoiser network.
-In the following, we detail the problem formulation of the 3D VM estimation using diffusion models, including the definition of the forward diffusion process and the reverse denoising process.
-Then, we introduce the architecture design of the denoiser network.
-Problem formulation.
-We formulate the 3D VM estimation as a reverse denoising process [26], as shown in the inference workflow at the bottom of Figure 3.
-Starting with a Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$, we design a denoiser network to iteratively estimate the clean 3D VM.
-To achieve this, the framework includes two core processes: a forward diffusion process and a reverse denoising process.
-(1) The forward diffusion process disturbs a clean 3D VM, i.e.
-the ground-truth (GT) 3D VM $\hat{\mathbf{P}}$ from the data distribution, by incrementally adding noise to generate a series of noisy 3D VMs and ultimately reaches a Gaussian prior.
-(2) Then the reverse denoising process iteratively denoises the Gaussian noise into clean data in a step-by-step manner.
-To guide the denoising process, we condition it on the 2D image cues $\mathbf{F}$ and the estimated 2D VMs $\mathbf{P}^{2d}\in\mathbb{R}^{2\times K}$, which is the first two dimensions of the estimated 3D VMs $\mathbf{P}$ by VMarker.
-By formulating the estimation as a reverse denoising process, we leverage the generative capabilities of diffusion models, framing the task as a probabilistic estimation naturally.
-Forward diffusion process.
-Starting from a clean 3D VM $\hat{\mathbf{P}}$ drawn from the GT data distribution, also denoted as $\hat{\mathbf{P}}(0)$, we construct a time-dependent diffusion process by incrementally adding noise to $\hat{\mathbf{P}}(0)$ to generate a series of noisy 3D VMs $\{\hat{\mathbf{P}}(t)\}_{t=0}^{T}$, indexed by a discrete timestep $t\in\{0,1,...,T\}$, where $T$ denotes the total number of timesteps.
-Formally, we define a noisy 3D VM $\hat{\mathbf{P}}(t)$ at timestep $t$ as:
-$q(\hat{\mathbf{P}}(1:T)|\hat{\mathbf{P}}(0))$ $:=\prod_{t=1}^{T}q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(t-1)),$ (11)
-$q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(t-1))$ $:=\mathcal{N}(\hat{\mathbf{P}}(t);\sqrt{1-\beta_{t}}\hat{\mathbf{P}}(t-1),\beta_{t}\mathbf{I}),$ (12)
-where $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule [26].
-Leveraging the additive property of independent Gaussian distributions and reparameterization techniques [75], we can obtain the perturbation $\hat{\mathbf{P}}(t)$ from a clean 3D VM $\hat{\mathbf{P}}(0)$ directly:
-$q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(0)):=\sqrt{\overline{\alpha}_{t}}\hat{\mathbf{P}}(0)+\sqrt{1-\overline{\alpha}_{t}}\hat{\bm{\epsilon}},$ (13)
-where $\alpha_{t}:=1-\beta_{t}$, $\overline{\alpha_{t}}:=\prod_{u=1}^{t}\alpha_{u}$, and $\hat{\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ as a Gaussian noise.
-Following previous work [26], we disturb any clean 3D VM $\hat{\mathbf{P}}(0)$ to a Gaussian distribution $\hat{\mathbf{P}}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
-Reverse denoising process.
-Starting from a Gaussian distribution, denoted as $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, we reverse the diffusion process by incrementally denoising it to get a clean 3D VM $\mathbf{P}(0)$.
-We adopt Denoising Diffusion Implicit Models (DDIM) [60] to accelerate the reverse denoising process with fewer timesteps while keeping the denoising quality.
-Formally, following DDIM [60], we define a timestep subset $\tau\subset\{0,1,...,T\}$, and the reverse process is defined as:
-$\begin{split}\mathbf{P}({\tau_{i-1}})&=\sqrt{\overline{\alpha}_{\tau_{i-1}}}(\frac{\mathbf{P}({\tau_{i}})-\sqrt{1-\overline{\alpha}_{\tau_{i}}}\bm{\epsilon}_{\tau_{i}}}{\sqrt{\overline{\alpha}_{\tau_{i}}}})\\ &+\sqrt{1-\overline{\alpha}_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\bm{\epsilon}_{\tau_{i}}+\sigma_{\tau_{i}}\hat{\bm{\epsilon}}_{\tau_{i}},\end{split}$ (14)
-where $\tau_{i},\tau_{i-1}$ are the adjacent timesteps in the subset $\tau$.
-$\sigma_{\tau_{i}}(\eta)=\eta\sqrt{(1-\overline{\alpha}_{\tau_{i-1}})/({1-\overline{\alpha}_{\tau_{i}}})}\sqrt{1-\overline{\alpha}_{\tau_{i}}/\overline{\alpha}_{\tau_{i-1}}}$, where $\eta\in\mathbb{R}_{\geq 0}$ is a hyperparameter to control the generation process.
-$\hat{\bm{\epsilon}}_{\tau_{i}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is standard Gaussian noise.
-In order to simulate Eq.
-14 to perform the denoising process, we need to know $\bm{\epsilon}_{\tau_{i}}$ for all timestep $\tau_{i}$.
-Therefore, we train a time-dependent denoiser network to estimate it as shown at the bottom block of Figure 3.
-Instead of directly predicting $\bm{\epsilon}_{\tau_{i}}$ as the original formulation by Ho et al.
-[26], following previous work [76, 68], we predict the clean signal itself, i.e.
-$\mathbf{P}(0)$, and obtain the noise $\bm{\epsilon}_{\tau_{i}}$ through:
-$\bm{\epsilon}_{\tau_{i}}=\frac{\mathbf{P}({\tau_{i}})-\sqrt{\overline{\alpha}_{{\tau_{i}}}}\mathbf{P}(0)}{\sqrt{1-\overline{\alpha}_{\tau_{i}}}}.$ (15)
-By combining Eq. 15 and Eq. 14, we can get the denoised $\mathbf{P}({\tau_{i-1}})$ as follows:
-$\mathbf{P}({\tau_{i-1}})$ $=\sqrt{\overline{\alpha}_{\tau_{i-1}}}\mathbf{P}(0)$ (16)
-$+\sqrt{1-\overline{\alpha}_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\frac{\mathbf{P}({\tau_{i}})-\sqrt{\overline{\alpha}_{{\tau_{i}}}}\mathbf{P}(0)}{\sqrt{1-\overline{\alpha}_{\tau_{i}}}}$
-$+\sigma_{\tau_{i}}\hat{\bm{\epsilon}}_{\tau_{i}}.$
-Denoiser architecture design.
-Figure 4 presents an overview of the time-dependent denoiser network architecture.
-The denoiser aims to estimate the clean 3D VM $\mathbf{P}(0)$ given a noisy 3D VM $\hat{\mathbf{P}}(t)$ at timestep $t$ as input.
-To guide the prediction, we concatenate $\hat{\mathbf{P}}(t)$ with the estimated 2D VM $\mathbf{P}^{2d}$, and then transform them to feature $\mathbf{F}^{vm}\in\mathbb{R}^{K\times C}$ by a multilayer perceptron (MLP), where $C$ denotes the number of feature channel.
-We then deploy a network with $B$ blocks to process the VM features $\mathbf{F}^{vm}$.
-Each block is composed of two Locally Connected Network (LCN) units [77, 78] sequentially.
-Within each unit, an LCN layer [77, 78] is followed by Group Normalization (GN) [79], a SiLU activation function [80], and Dropout [81] for regularization.
-The LCN layer, as proposed in [77, 78], is akin to an enhanced version of MLP, offering more effective feature processing across virtual markers.
-To improve the alignment of the denoised 3D VM with 2D image cues, we extract 2D image features $\mathbf{F}\in\mathbb{R}^{H_{b}\times W_{b}\times C_{b}}$ using the CNN backbone [18].
-These features are sampled according to the estimated 2D VMs $\mathbf{P}^{2d}$ and subsequently transformed into a new feature space $\mathbf{F}^{2d}\in\mathbb{R}^{K\times C}$ through a linear layer, as illustrated in Figure 4.
-We then fuse $\mathbf{F}^{2d}$ into the network by Cross-Attention [82].
-To achieve this, we slightly modify the LCN layer by incorporating an attention module [82].
-In addition, we process the current timestep $t$ using an MLP to be $\mathbf{F}^{t}\in\mathbb{R}^{1\times C}$ and add it to the processed VM features $\mathbf{F}^{vm}$ to make the denoiser time-dependent following common practice [74, 53].
-Finally, an FC layer is deployed to estimate the clean signal $\mathbf{P}(0)$.
-Please refer to the supplementary for more details.
-3.3.2 Training
-The whole framework is trained in an end-to-end manner. The overall loss function is defined as:
-$\mathcal{L}_{vmpro}=\mathcal{L}_{vm}+\lambda_{diffusion}\mathcal{L}_{diffusion},$ (17)
-where $\mathcal{L}_{vm}$ is the loss function for VMarker, as defined in Eq.
-4.
-Following common practice in training diffusion models [26, 68], we randomly sample a timestep $t\in\{0,1,...,T\}$ to get the noisy 3D VM $\hat{\mathbf{P}}(t)$ according to Eq.
-13 and deploy $\mathcal{L}_{diffusion}$ as the $L_{1}$ distance between the predicted clean 3D VM $\mathbf{P}(0)$ and the GT $\hat{\mathbf{P}}(0)$:
-$\mathcal{L}_{diffusion}=\mathbb{E}_{t,\hat{\mathbf{P}}(0),\hat{\bm{\epsilon}}}[\|\mathbf{P}(0)-\hat{\mathbf{P}}(0)\|_{1}].$ (18)
-3.3.3 Inference
-As shown in Figure 3, at inference, the denoiser iteratively removes noise from a Gaussian noise $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ using DDIM [60] according to Eq.
-16, yielding a denoised 3D VM $\mathbf{P}(0)$.
-Finally, we obtain a mesh estimate $\mathbf{M}=\mathbf{P}(0)\mathbf{A}$.
-By sampling $S$ Gaussian noises, we can obtain $S$ feasible mesh estimates.
-Figure 4: Architecture of the denoiser network, which predicts the clean 3D VM $\mathbf{P}(0)$ from the noisy 3D VM $\hat{\mathbf{P}}(t)$.
-To guide the direction of denoising, we additionally input 2D features $\mathbf{F}$, the current timestep $t$, and the 2D VM $\mathbf{P}^{2d}$ predicted by VMarker.
-We concatenate $\mathbf{P}^{2d}$ with the noisy 3D VM $\hat{\mathbf{P}}(t)$ to facilitate training and fuse the sampled 2D features $\mathbf{F}$ through a Cross-Attention mechanism [82].
-We utilize a network comprising $B$ blocks to process the features, where each block contains two Locally Connected Network (LCN) units [77, 78].
-Each LCN units consists of one LCN layer [77, 78], followed by Group Normalization (GN) [79], SiLU activation [80], and Dropout [81].
-Finally, an FC layer is used to predict the clean 3D VM $\mathbf{P}(0)$.
+    $\mathcal{L}_{normal}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\left<\frac{\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}}{\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}},\hat{\mathbf{n}}_{f}^{*}\right>\right|,$ (10)
+    $\mathcal{L}_{edge}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}-\|\hat{\mathbf{M}}_{i}^{*}-\hat{\mathbf{M}}_{j}^{*}\|_{2}\right|.$
+    where $f$ and $\hat{\mathbf{n}}_{f}^{*}$ denote a triangle face in the mesh and its GT unit normal vector, respectively.
+$\hat{\mathbf{M}}_{i}$ denote the $i^{th}$ vertex of $\hat{\mathbf{M}}$.
+denotes GT.
 4 Experiments
-TABLE II: Comparison to the deterministic state-of-the-arts on H3.6M [13] and 3DPW [14] datasets.
-means using temporal cues.
-The methods are not strictly comparable because they may have different backbones and training datasets.
-We provide the numbers only to show proof-of-concept results.
-Method Venue Intermediate H3.6M [13] 3DPW [14]
+Method Intermediate H3.6M 3DPW
 […]
+Arnab et al. [1] CVPR’19 2D skeleton - 77.8 54.3 - - 72.2
 HMMR [20] CVPR’19 - - - 56.9 139.3 116.5 72.6
@@
 PyMaf [64] ICCV’21 IUV image - 57.7 40.5 110.1 92.8 58.9
+ProHMR [27] ICCV’21 - - - 41.2 - - 59.8
 OCHMR [21] CVPR’22 2D heatmap - - - 107.1 89.7 58.3
@@
 VisDB [58] ECCV’22 3D vertices - 51.0 34.5 85.5 73.5 44.9
-VMarker (Ours) [28] CVPR’23 Virtual marker 58.0 47.3 32.0 77.9 67.5 41.3
-TABLE III: Comparison to the probabilistic state-of-the-arts on H3.6M [13] and 3DPW [14] datasets.
+Ours Virtual marker 58.0 47.3 32.0 77.9 67.5 41.3
+Table 2: Comparison to the state-of-the-arts on H3.6M [16] and 3DPW [54] datasets.
+means using temporal cues.
 The methods are not strictly comparable because they may have different backbones and training datasets.
 We provide the numbers only to show proof-of-concept results.
-Method Hypothesis H3.6M [13] 3DPW [14]
-[…]
-Biggs et al.  [41] NeurIPS’20 10 - 59.2 42.2 - 79.4 56.6
-25 - 58.2 42.2 - 75.8 55.6
-Sengupta et al.  [56] ICCV’21 25 - - - - 75.1 47.0
-ProHMR [44] ICCV’21 10 - - - - 88.9 55.0
-25 - - - - 85.1 52.1
-100 - - - - 80.1 48.1
-200 - - - - 77.9 46.5
-HuManiFlow [43] CVPR’23 10 - - - - 75.6 47.9
-25 - - - - 71.9 44.5
-100 - - - - 65.1 39.9
-200 - - - - 64.5 38.8
-HMDiff [45] ICCV’23 25 - 49.3 32.4 82.4 72.7 44.5
-10 55.8 45.4 31.3 77.3 66.6 42.5
-25 52.9 42.9 29.8 72.9 63.2 40.0
-100 49.2 39.9 27.6 67.4 58.2 36.4
-VMarker-Pro (Ours) 200 47.7 38.6 26.6 64.7 56.2 35.0
 4.1 Datasets and metrics
@@
 We use (S1, S5, S6, S7, S8) for training and (S9, S11) for testing.
-For the deterministic estimation task, as in [6, 1, 87, 89], we report Mean Per Joint Position Error (MPJPE) and PA-MPJPE for poses that are derived from the estimated meshes.
+As in [19, 7, 32, 33], we report MPJPE and PA-MPJPE for poses that are derived from the estimated meshes.
 We also report Mean Per Vertex Error (MPVE) for the whole mesh.
-For the probabilistic estimation task, we report the minimum errors over $S$ hypotheses to measure the estimated distribution accuracy following a standard practice in recent works [43, 45].
 3DPW [54] is collected in natural scenes.
-Following previous works [87, 89, 47, 40, 43, 45], we use the train set of 3DPW to learn the model and evaluate on the test set.
+Following the previous works [32, 33, 24, 61], we use the train set of 3DPW to learn the model and evaluate on the test set.
 The same evaluation metrics as H3.6M are used.
-We further evaluate the performance of our probabilistic framework VMarker-Pro on two subsets of 3DPW, i.e.
-3DPW-OC [14, 95] and 3DPW-PC [14, 88] which are composed by object- and person-specific occlusion, respectively.
-To ensure a fair comparison, as suggested by [96], we do not use the 3DPW training set in these particular evaluations.
 SURREAL [53] is a large-scale synthetic dataset with GT SMPL annotations and has diverse samples in terms of body shapes, backgrounds, etc.
 We use its training set to train a model and evaluate the test split following [7].
-The same evaluation metrics as H3.6M [13] are reported.
 4.2 Implementation Details
 We learn $64$ virtual markers on the H3.6M [16] training set.
-We use the same set of markers for all datasets instead of learning a separate set for each one.
+We use the same set of markers for all datasets instead of learning a separate set on each dataset.
 Following [19, 7, 39, 61, 26, 23, 33, 32], we conduct mix-training by using MPI-INF-3DHP [38], UP-3D [28], and COCO [34] training set for experiments on the H3.6M and 3DPW datasets.
@@
 Following [19, 26, 39], we crop every single human region from the input image and resize it to $256\times 256$.
-The dimensions of the 2D feature map are $64$, i.e.
-$H_{b}=W_{b}=64$.
-The 2D feature map has $C_{b}=48$ feature channels.
-The processed $\mathbf{F}^{2d}$ feature channels are set to be $C=64$.
-The denoiser network has $B=3$ blocks.
-To train VMarker, we use Adam [100] optimizer to train the whole framework for $40$ epochs with a batch size of $32$.
-The learning rates for the 3D VM estimation branch and the updating branch are set to $5\times 10^{-4}$ and $1\times 10^{-3}$, respectively, which are decreased by half after the $30^{th}$ epoch.
-To train VMarker-Pro, we use Adam [100] optimizer to train the whole framework for $50$ epochs with a batch size of $80$.
-The learning rates for the 3D VM estimation branch, the updating branch, and the denoiser are set to $4\times 10^{-5}$, $2\times 10^{-4}$, and $2\times 10^{-3}$, respectively.
-They are decayed by $0.5$, $0.5$, and $0.1$ after $30^{th}$ and $40^{th}$ epochs, respectively.
-For inference, we employ DDIM [60] strategy and generate each hypothesis in $T=10$ steps with $\eta=0$.
-Please check the supplementary for more details.
-TABLE IV: Comparison to the state-of-the-arts on SURREAL [27] dataset.
-means training on the test split with 2D supervisions.
-“Skel.
-+ Seg.” means using skeleton and segmentation together.
-$S$ is the number of hypotheses.
+We use Adam [22] optimizer to train the whole framework for $40$ epochs with a batch size of $32$.
+The learning rates for the two branches are set to $5\times 10^{-4}$ and $1\times 10^{-3}$, respectively, which are decreased by half after the $30^{th}$ epoch.
+Please refer to the supplementary for more details.
 […]
 Representation
-HMR [6] - - 85.1 73.6 55.4
-BodyNet [46] Skel. + Seg. - 65.8 - -
-GraphCMR [16] 3D vertices - 103.2 87.4 63.2
-SPIN [33] - - 82.3 66.7 43.7
-DecoMR [38] IUV image - 68.9 52.0 43.0
-Pose2Mesh [1] 3D skeleton - 68.8 56.6 39.6
-PC-HMR [36] 3D skeleton - 59.8 51.7 37.9
-DynaBOA [101] - - 70.7 55.2 34.0
-VMarker (Ours) Virtual marker - 44.7 36.9 28.9
-10 44.3 36.6 28.0
-VMarker-Pro 25 42.0 34.8 26.2
-(Ours) 100 38.1 31.0 23.5
-Virtual marker 200 36.5 29.8 22.4
-TABLE V: Comparison to the state-of-the-art methods on the challenging occlusion-specific 3DPW-OC [14, 95] and 3DPW-PC [14, 88] datasets.
-The top and bottom blocks are deterministic and probabilistic methods, respectively.
-Method Venue Hypothesis 3DPW-OC [14, 95] 3DPW-PC [14, 88]
-[…]
-SPIN [33] ICCV’19 - 121.4 95.5 60.7 159.8 122.1 77.4
-I2L-MeshNet [17] ECCV’20 - 129.5 92.0 61.4 160.2 117.3 80.0
-PyMAF [37] ICCV’21 - 113.7 89.6 59.1 154.6 117.5 74.5
-ROMP [88] ICCV’21 - - 91.0 62.0 152.8 117.9 79.7
-PARE [47] ICCV’21 - 101.5 83.5 57.0 122.4 96.8 64.5
-OCHMR [90] CVPR’22 - 145.9 112.2 75.2 149.6 117.5 77.1
-3DCrowdNet [91] CVPR’22 - 101.5 83.5 57.1 114.8 90.9 64.4
-JOTR [96] ICCV’23 - 92.6 75.7 52.2 109.7 86.5 58.3
-VMarker (Ours) [28] CVPR’23 - 92.8 78.4 49.0 110.8 93.9 66.0
-HMDiff [45] ICCV’23 25 - - - 143.1 114.2 73.5
-10 90.4 77.3 48.6 109.4 89.8 60.2
-25 85.4 73.3 46.0 102.9 84.7 56.6
-100 77.7 67.1 42.5 94.0 78.0 52.0
-VMarker-Pro (Ours) - 200 75.2 64.9 40.9 90.3 74.6 49.7
+HMR [19] CVPR’18 - 85.1 73.6 55.4
+BodyNet [52] ECCV’18 Skel. + Seg. 65.8 - -
+GraphCMR [26] CVPR’19 3D vertices 103.2 87.4 63.2
+SPIN [25] ICCV’19 - 82.3 66.7 43.7
+DecoMR [62] CVPR’20 IUV image 68.9 52.0 43.0
+Pose2Mesh [7] ECCV’20 3D skeleton 68.8 56.6 39.6
+PC-HMR [37] AAAI’21 3D skeleton 59.8 51.7 37.9
+DynaBOA [14] TPAMI’22 - 70.7 55.2 34.0
+Ours Virtual marker 44.7 36.9 28.9
+Table 3: Comparison to the state-of-the-arts on SURREAL [53] dataset. means training on the test split with 2D supervisions. “Skel. + Seg.” means using skeleton and segmentation together.
 4.3 Comparison to the State-of-the-arts
-4.3.1 Deterministic Track
-Results on H3.6M [13].
-Table II compares our approach to the deterministic state-of-the-art (SOTA) methods on the H3.6M dataset.
-Our method achieves superior performance.
+Results on H3.6M.
+Table 2 compares our approach to the state-of-the-art methods on the H3.6M dataset.
+Our method achieves competitive or superior performance.
 In particular, it outperforms the methods that use skeletons (Pose2Mesh [7], DSD-SATN [49]), body markers (THUNDR) [61], or IUV image [62, 64] as proxy representations, demonstrating the effectiveness of the virtual marker representation.
-Results on 3DPW [14].
-We compare our method to the deterministic SOTA methods on the 3DPW dataset in Table II.
-Our approach achieves SOTA results among all the methods, validating the advantages of the virtual marker representation over the skeleton representation used in Pose2Mesh [1], DSD-SATN [48], and other representations like IUV image used in PyMAF [37].
-In particular, our approach outperforms I2L-MeshNet [17], METRO [87], and Mesh Graphormer [89] by a notable margin, suggesting that virtual markers are more suitable and effective representations than using all vertices as most of them are not discriminative enough to be accurately detected.
-Results on SURREAL [27].
+Results on 3DPW.
+We compare our method to the state-of-the-art methods on the 3DPW dataset in Table 2.
+Our approach achieves state-of-the-art results among all the methods, validating the advantages of the virtual marker representation over the skeleton representation used in Pose2Mesh [7], DSD-SATN [49], and other representations like IUV image used in PyMAF [64].
+In particular, our approach outperforms I2L-MeshNet [39], METRO [32], and Mesh Graphormer [33] by a notable margin, which suggests that virtual markers are more suitable and effective representations than detecting all vertices directly as most of them are not discriminative enough to be accurately detected.
+Results on SURREAL.
 This dataset has more diverse samples in terms of body shapes.
-The results are shown in Table IV.
-Our approach VMarker, outperforms the SOTA methods by a notable margin, especially in terms of MPVE.
-Figure 1 (top) shows some challenging cases without cherry-picking.
+The results are shown in Table 3.
+Our approach outperforms the state-of-the-art methods by a notable margin, especially in terms of MPVE.
+Figure 1 shows some challenging cases without cherry-picking.
 The skeleton representation loses the body shape information so the method [7] can only recover mean shapes.
 In contrast, our approach generates much more accurate mesh estimation results.
-4.3.2 Probabilistic Track
-Results on H3.6M [13] and 3DPW [14].
-Table III presents a comparison between VMarker-Pro and other SOTA probabilistic methods on the H3.6M [13] and 3DPW [14] datasets.
-Our method significantly surpasses the existing SOTA approaches in terms of accuracy when estimating data distributions.
-This serves as a validation of the effectiveness of our probabilistic framework, which benefits from the generation capability of diffusion models.
-Furthermore, as the number of hypotheses ($S$) increases, our method consistently improves, highlighting the diverse and accurate nature of our approach when modeling data distributions.
-Notably, compared to HMDiff [45], which is also a diffusion model-based method, our incorporation of 2D image cues and the virtual marker representations contributes to the observed notable improvement.
-Results on 3DPW-OC [14, 95] and 3DPW-PC [14, 88].
-Table V shows the comparison results of our method against existing methods on the 3DPW-OC and 3DPW-PC datasets.
-Both 3DPW-OC [14, 95] and 3DPW-PC [14, 88] are subsets of the 3DPW [14] dataset, containing scenarios with occlusions by objects and people, respectively.
-It can be seen that our deterministic estimation results are comparable to the SOTA performance.
-When adopting the proposed probabilistic modeling approach, the accuracy of our method in estimating the data distribution is significantly improved, thanks to the enhanced accuracy in 3D virtual marker estimation.
-Figure 5 presents a comparative visualization between VMarker and VMarker-Pro.
-It can be seen that the deterministic approach, VMarker, may encounter difficulties under occlusions, e.g.
-the wrongly estimated right arm and the incorrect arm length.
-Conversely, the probabilistic method VMarker-Pro, demonstrates a capability to accurately model realistic human poses, offering feasible estimations for the occluded parts.
-Figure 5: Qualitative comparison of VMarker and VMarker-Pro on 3DPW-OC [14, 88] testset.
-For each mesh, we show a projecting view and a side view (with shadow).
-VMarker wrongly estimates the right arm when heavy occlusion happens while VMarker-Pro provides two reasonable estimates.
-Results on SURREAL [27].
-Table IV displays the results of our probabilistic approach on the SURREAL [27] dataset.
-It can be seen that when $S=10$, our probabilistic framework already surpasses existing approaches.
-This indicates that our method can accurately model the more diverse data distribution.
-As the number of hypotheses ($S$) increases, the performance of our model continues to improve.
-4.4 Ablation Study on VMarker
+[…]
+Representation H3.6M SURREAL
+(a) Skeleton 64.4 53.6
+(b) Rand virtual marker 63.0 50.1
+(c) Virtual marker 58.0 44.7
+Table 4: Ablation study of the virtual marker representation for our approach on H3.6M and SURREAL datasets.
+“Skeleton” means the sparse landmark joint representation is used.
+“Rand virtual marker” means the virtual markers are randomly selected from all the vertices without learning.
+(c) is our method, where the learned virtual markers are used.
+4.4 Ablation study
 Virtual marker representation.
-We compare our method to two baselines in Table VI.
+We compare our method to two baselines in Table 4.
 First, in baseline (a), we replace the virtual markers of our method with the skeleton representation.
@@
 We repeat the experiment five times and report the average number.
-We can see that the result is worse than ours, which is because the randomly selected vertices may not be expressive enough to reconstruct the other vertices or can not be accurately detected from images as they lack distinguishable visual patterns.
+We can see that the result is worse than ours, which is because the randomly selected vertices may not be expressive to reconstruct the other vertices or can not be accurately detected from images as they lack distinguishable visual patterns.
 The results validate the effectiveness of our learning strategy.
-TABLE VI: Ablation study of the virtual marker representation for our approach VMarker on H3.6M [13] and SURREAL [27] datasets.
-“Skeleton” means the sparse landmark joint representation is used.
-“Rand virtual marker” means the virtual markers are randomly selected from all the vertices without learning.
-(c) is our method, where the learned virtual markers are used.
-MPVE error is reported.
-[…]
-Representation H3.6M SURREAL
-(a) Skeleton 64.4 53.6
-(b) Rand virtual marker 63.0 50.1
-(c) Virtual marker 58.0 44.7
-Figure 6: Mesh estimation results of different methods on H3.6M [13] test set.
-Our method VMarker with virtual marker representation gets better shape estimation results than Pose2Mesh [1] which uses skeleton representation.
+Figure 4: Mesh estimation results of different methods on H3.6M test set.
+Our method with virtual marker representation gets better shape estimation results than Pose2Mesh which uses skeleton representation.
 Note the waistline of the body and the thickness of the arm.
-Figure 1 (top) shows some qualitative results on the SURREAL test set.
+Figure 1 shows some qualitative results on the SURREAL test set.
 The meshes estimated by the baseline which uses skeleton representation, i.e.
@@
 Figure 4 also shows some qualitative results on the H3.6M test set.
-For clarity, we also draw the intermediate representation (blue balls) in it.
+For clarity, we draw the intermediate representation (blue balls) in it as well.
+Figure 5: Visualization of the learned virtual markers of different numbers of $K=16,32,96$, from left to right, respectively.
 Number of virtual markers.
 We evaluate how the number of virtual markers affects estimation quality on H3.6M [16] dataset.
-Figure 2 visualizes the learned 64 markers.
-Figure 7 visualizes the learned virtual markers when $K=16,32,96$, which are all located on the body surface and close to the extreme points of the mesh.
+Figure 5 visualizes the learned virtual markers, which are all located on the body surface and close to the extreme points of the mesh.
 This is expected as mentioned in Section 3.1.
-Table VII (GT) shows the mesh reconstruction results when we have GT 3D positions of the virtual markers in objective (Eq.
-1).
-When we increase the number of virtual markers, both mesh reconstruction error, i.e.
-MPVE, and the regressed landmark joint error, i.e.
-MPJPE, steadily decrease.
+Table 5 (GT) shows the mesh reconstruction results when we have GT 3D positions of the virtual markers in objective (1).
+When we increase the number of virtual markers, both mesh reconstruction error (MPVE) and the regressed landmark joint error (MPJPE) steadily decrease.
 This is expected because using more virtual markers improves the representation power.
@@
 This is because the additional virtual markers may have large estimation errors which affect the mesh estimation result.
-The results are shown in Table VII (Est).
+The results are shown in Table 5 (Det).
 Increasing the number of virtual markers $K$ steadily reduces the MPVE errors when $K$ is smaller than $96$.
@@
 This is mainly because some of the newly introduced virtual markers are difficult to detect from images and therefore bring errors to mesh estimation.
-TABLE VII: Ablation study of the different number of virtual markers ($K$) on H3.6M [13] dataset.
-(GT) Mesh reconstruction results when GT 3D positions of the virtual markers are used in objective (1).
-(Est) Mesh estimation results obtained by our method VMarker when we use different numbers of virtual markers ($K$).
-$K$ GT Est
+$K$ GT Det
 […]
@@
 96 9.9 5.6 59.6 48.2
-Figure 7: Visualization of the learned virtual markers when $K=16,32,96$, from left to right, respectively.
-Updating coefficient matrix.
+Table 5: Ablation study of the different number of virtual markers ($K$) on H3.6M [16] dataset.
+(GT) Mesh reconstruction results when GT 3D positions of the virtual markers are used in objective (1).
+(Det) Mesh estimation results obtained by our proposed framework when we use different numbers of virtual markers ($K$).
+Figure 6: Mesh estimation comparison results when using (a) fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$, and (b) updated $\hat{\mathbf{A}}$. Please zoom in to better see the details.
+Figure 7: Top: Meshes estimated by our approach on images from 3DPW test set.
+The rightmost case in the dashed box shows a typical failure.
+Bottom: Meshes estimated by our approach on Internet images with challenging cases (extreme shapes or in a long dress).
+Coefficient matrix.
 We compare our method to a baseline which uses the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$.
 We show the quality comparison in Figure 6.
-We can see that the estimated mesh by (a) a fixed coefficient matrix has mostly correct pose and shape but there are some artifacts on the mesh while using the (b) updated coefficient matrix can get better mesh estimation results.
-As shown in Table VIII, using a fixed coefficient matrix gets larger MPVE and MPJPE errors than using the updated one.
+We can see that the estimated mesh by a fixed coefficient matrix (a) has mostly correct pose and shape but there are also some artifacts on the mesh while using the updated coefficient matrix (b) can get better mesh estimation results.
+As shown in Table 6, using a fixed coefficient matrix gets larger MPVE and MPJPE errors than using the updated coefficient matrix.
 This is caused by the estimation errors of virtual markers when occlusion happens, which is inevitable since the virtual markers on the back will be self-occluded by the front body.
 As a result, inaccurate marker positions would bring large errors to the final mesh estimates if we directly use the fixed matrix.
-TABLE VIII: Ablation study of the coefficient matrix for VMarker on H3.6M [13] dataset. “fixed” means using the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ to reconstruct the mesh.
 […]
-(a) VMarker (fixed) 64.7 51.6
-(b) VMarker 58.0 47.3
-Figure 8: Mesh estimation comparison results when using (a) fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$, and (b) updated $\mathbf{A}$ in our method VMarker.
-Please zoom in to better see the details.
-TABLE IX: Ablation study of the architecture design for the probabilistic method VMarker-Pro on 3DPW [14] dataset.
-(a-b) ablates the use of 2D feature $\mathbf{F}$ and 2D VMs $\mathbf{P}^{2d}$ in denoiser network, respectively.
-Ablation (c) removes the 2D VM $\mathbf{P}^{2d}$ sampling in 2D feature $\mathbf{F}$ and directly transforms the global 2D feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ for the denoiser.
-In ablation (d), the denoiser directly estimates the noise $\bm{\epsilon}$ following the original formula [26], while keeping the rest design the same.
-(e) is the proposed VMarker-Pro framework.
-[…]
-$\mathbf{F}$ $\mathbf{P}^{2d}$ $\mathbf{H}^{2d}$
-(a) w/o $\mathbf{F}$ 80.1 68.2
-(b) w/o $\mathbf{P}^{2d}$ 80.6 68.8
-(c) w/o sampling in $\mathbf{F}$ 80.4 69.2
-(d) $\rightarrow\bm{\epsilon}$ 80.6 68.9
-(e) VMarker-Pro 76.9 66.1
-Figure 9: Mesh estimation comparison results of different ablated methods of VMarker-Pro.
-(a) We ablate the usage of 2D image features $\mathbf{F}$.
-(b) We ablate the usage of 2D estimated VM $\mathbf{P}^{2d}$.
-(c) We remove the 2D VM $\mathbf{P}^{2d}$ sampling in 2D feature $\mathbf{F}$ and directly transform the global 2D feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ for the denoiser.
-(d) We change the estimation target of the denoiser to the noise $\bm{\epsilon}$ instead of the signal itself.
-(e) is our proposed VMarker-Pro method.
-We show the GT mesh and its side view on the far right.
-For baselines (a-e), we show each hypothesis that is denoised from a zero noise for a fair comparison.
-Each hypothesis presents a projecting view and a side view.
-4.5 Ablation Study on VMarker-Pro
-Architecture design of the denoiser.
-Table IX compares VMarker-Pro to four baselines.
-We report the metric when predicting $S=10$ hypotheses.
-In baseline (a) and (b), we remove the 2D feature $\mathbf{F}$ and 2D estimated VMs $\mathbf{P}^{2d}$ input for the denoiser, respectively.
-It can be seen that the errors increase a lot compared to the full model, which validates the effectiveness of the two guidance.
-In baseline (c), we remove the local sampling by 2D VMs $\mathbf{P}^{2d}$ on 2D feature $\mathbf{F}$, i.e.
-the purple arrow in Figure 4.
-Instead, we directly transform the global 2D image feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ when feeding into the denoiser and the performance degrades notably.
-In ablation study (d), we alter the denoiser to estimate the noise $\bm{\epsilon}$ rather than the clean 3D VMs $\mathbf{P}(0)$, adjusting the inference process to align with the original denoising formulation by Ho et al.
-[26].
-This modification results in a decrease in performance.
-A similar observation was made in previous research by Tevet et al.
-[68], where it is suggested that noise estimation might require substantially larger batch sizes to be effective.
-Unfortunately, we lack the resources to confirm this hypothesis in our settings.
-Figure 9 presents the qualitative comparison of different designs.
-As observed in the image, the left arm of the subject is nearly fully extended.
-However, due to depth ambiguity, the estimations from baselines (a-d) align with the 2D perspective, yet inaccuracies arise in the 3D estimation of the left forearm when viewed from the side view.
-Our method (e), which integrates 2D local features, more accurately captures precise 3D poses, achieving estimations that most closely match the GT.
-This highlights the significance of our design, which not only utilizes 2D local features $\mathbf{F}$ but also leverages 2D VMs $\mathbf{P}^{2d}$ to locally sample these features effectively.
-TABLE X: Results of using different denoising steps $T$ for VMarker-Pro on H3.6M [13] and 3DPW [14] datasets.
-[…]
-[…]
-1 386.5 370.4 372.2 358.9 339.6
-2 77.0 66.2 104.7 92.6 326.6
-4 58.8 48.3 80.0 69.2 301.0
-5 57.1 46.9 78.3 67.6 292.6
-10 55.8 45.4 77.3 66.6 247.4
-15 55.5 45.0 76.4 65.5 208.0
-20 55.8 45.3 76.8 65.9 189.0
-100 55.6 45.2 76.9 65.9 66.4
-Figure 10: Mesh estimation results of using different denoising steps $T$ for VMarker-Pro on a sample data from H3.6M [13] dataset.
-Each hypothesis is denoised from a zero initial noise for a fair comparison.
-Denoising steps.
-Table X shows the metrics on the H3.6M [13] and 3DPW [14] test sets when using different numbers of denoising steps $T$.
-We report the metrics of predicting $S=10$ hypotheses.
-We also report the speed of the whole model inferring a hypothesis on an NVIDIA A100-PCIE-80GB GPU.
-It can be seen that using only $T=1$ denoising step results in significant errors, and the estimated mesh is not human-shaped, as shown in Figure 10.
-As the number of denoising steps $T$ increases, the accuracy of the estimated mesh gradually improves, and the error decreases.
-When $T>=10$, the accuracy of the mesh estimation stabilizes, and further steps yield only marginal improvements, as validated by the marginal reduction in error and the indistinguishable differences in visualization.
-Therefore, based on the acceleration of DDIM [60], we choose $T=10$ steps for a good balance between estimation performance and inference speed.
-TABLE XI: Robustness to occlusion for VMarker and VMarker-Pro on 3DPW [14] test set when different body parts are occluded. We report MPVE error.
-Occ. Parts VMarker VMarker-Pro
-$S=10$ $S=100$
-None 77.9 77.3 67.4
-2 Arms 79.2 78.7 68.8
-2 Legs 78.3 77.6 67.8
-Body 78.6 78.1 68.2
-Random 78.7 77.9 68.4
-Figure 11: Top: Meshes estimated by our deterministic approach VMarker on images from 3DPW test set.
-The rightmost case in the dashed box shows a typical failure.
-Bottom: Meshes estimated by VMarker on Internet images with challenging cases (extreme shapes or in a long dress).
-Figure 12: (a) For each virtual marker (represented by a star), we highlight the top 30 most affected vertices (represented by a colored dot) based on average coefficient matrix $\mathbf{A}$.
-(b) For each vertex (dot), we highlight the top 3 virtual markers (star) that contribute the most.
-The dependency has a strong locality which improves the robustness when some virtual markers cannot be accurately detected.
-Figure 13: Qualitative comparison of VMarker and VMarker-Pro on 3DPW [14] testset.
-VMarker may fail when facing ambiguity, while VMarker-Pro could generate multiple reasonable solutions.
-For each mesh estimate, we present both a projected view and a side view in each column.
-We show three hypotheses generated by VMarker-Pro.
-Figure 14: Meshes estimated by VMarker-Pro on Internet images.
-For each case in a row, we present three hypotheses, showing both a projected view and a side view in each column.
-The rightmost column shows the overlapped side views to unveil the subtle differences.
-Figure 15: A typical failure case of VMarker-Pro. The heavy occlusion and ambiguity cause the model to incorrectly estimate the right arm and right leg of the male dancer.
-Robustness to occlusion.
-We report results of VMarker and VMarker-Pro when different image regions of corresponding virtual markers are occluded by a synthetic random mask in Table XI.
-The errors of VMarker are slightly larger than the original image (None).
+(a) Ours (fixed) 64.7 51.6
+(b) Ours 58.0 47.3
+Table 6: Ablation study of the coefficient matrix for our approach on H3.6M dataset. “fixed” means using the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ to reconstruct the mesh.
+4.5 Qualitative Results
+Figure 7 (top) presents some meshes estimated by our approach on natural images from the 3DPW test set.
+The rightmost case shows a typical failure where our method has a wrong pose estimate of the left leg due to heavy occlusion.
+We can see that the failure is constrained to the local region and the rest of the body still gets accurate estimates.
 We further analyze how inaccurate virtual markers would affect the mesh estimation, i.e.
 when part of human body is occluded or truncated.
-According to the finally learned coefficient matrix $\mathbf{A}$ of our model VMarker, we highlight the relationship weights among virtual markers and all vertices in Figure 12.
+According to the finally learned coefficient matrix $\mathbf{\hat{A}}$ of our model, we highlight the relationship weights among virtual markers and all vertices in Figure 8.
 We can see that our model actually learns local and sparse dependency between each vertex and the virtual markers, e.g.
@@
 Therefore, in inference, if a virtual marker has inaccurate position estimation due to occlusion or truncation, the dependent vertices may have inaccurate estimates, while the rest will be barely affected.
-This locality enhances the robustness of the VMarker-Pro method against occlusion as well.
-When sampling $S=10$ hypotheses, VMarker-Pro exhibits a minor increase in error under various circumstances of VM occlusion.
-4.6 Qualitative Results
-VMarker.
-Figure 11 (top) presents some meshes estimated by our approach VMarker on natural images from the 3DPW test set.
-The rightmost case shows a typical failure where our method has a wrong pose estimate of the left arm due to heavy occlusion.
-We can see that the failure is constrained to the local region and the rest of the body still gets accurate estimates.
-As discussed in Sec.
-15, we suppose that this is due to the advantages brought by the locality of the virtual marker representation, which can confine erroneous mesh estimates to a local range and thereby enhance robustness.
 Figure 2 (right) shows more examples where occlusion or truncation occurs, and our method can still get accurate or reasonable estimates robustly.
 Note that when truncation occurs, our method still guesses the positions of the truncated virtual markers.
+Figure 8: (a) For each virtual marker (represented by a star), we highlight the top 30 most affected vertices (represented by a colored dot) based on average coefficient matrix $\mathbf{\hat{A}}$.
+(b) For each vertex (dot), we highlight the top 3 virtual markers (star) that contribute the most.
+We can see that the dependency has a strong locality which improves the robustness when some virtual markers cannot be accurately detected.
 Figure 7 (bottom) shows our estimated meshes on challenging cases, which indicates the strong generalization ability of our model on diverse postures and actions in natural scenes.
@@
 Note that since the datasets do not provide supervision of head orientation, face expression, hands, or feet, the estimates of these parts are just in canonical poses inevitably.
-VMarker-Pro.
-Figure 13 compares the visualization of VMarker and VMarker-Pro on the 3DPW [14] test set.
-We draw three hypotheses generated by VMarker-Pro, depicted in blue.
-Additionally, we include the results from the deterministic model VMarker (green).
-It is noticeable that, due to self-occlusion or object occlusion, VMarker might yield results that seem accurate in 2D projections yet are erroneous in 3D space, e.g.
-the incorrect estimation of the right leg in the second case.
-This issue stems from the inherent ambiguity associated with occlusions.
-By leveraging the advantages of the data distribution learned by the denoising model, VMarker-Pro can generate more reasonable solutions, such as the naturally standing legs.
-In Figure 14, we show more results of VMarker-Pro on unseen Internet images.
-It can be observed that VMarker-Pro possesses strong generalization capabilities, obtaining reasonable 3D estimations when faced with uncertain ambiguities, such as the various plausible poses of the runner’s right arm in the second case (as can be seen from the side view).
-Furthermore, even under complex actions and backgrounds, such as the person climbing in the third image, VMarker-Pro also demonstrates robust estimation results.
-Figure 15 shows a typical failure case of VMarker-Pro on Internet image, where our method fails when heavy occlusion and ambiguity happen.
-Our method incorrectly estimates the right arm and right leg of the male dancer, which may be addressed using more powerful estimators or more diverse training datasets in the future.
+Apart from that, most errors are due to inaccurate 3D virtual marker estimation which may be addressed using more powerful estimators or more diverse training datasets in the future.
 5 Conclusion
 In this paper, we present a novel intermediate representation Virtual Marker, which is more expressive than the prevailing skeleton representation and more accessible than physical markers.
-Our VMarker method, built on virtual markers, excels in detecting these markers in wild images to interpolate and reconstruct complete, realistically shaped meshes.
-To further address the ambiguity in 3D virtual marker prediction due to occlusions, we progress to VMarker-Pro, a probabilistic approach that generates multiple plausible meshes in alignment with the image.
-VMarker-Pro effectively captures the 3D virtual marker distribution through a conditional diffusion process, enabling precise modeling of 3D mesh distributions.
-Experiments demonstrate the effectiveness of our approaches and they show advanced generalization potential in spite of its simplicity.
-Acknowledgments
-This work was supported by National Science and Technology Major Project (2022ZD0114904).
-Appendix A Post-processing on Virtual Markers
+It can reconstruct 3D meshes more accurately and efficiently, especially in handling diverse body shapes.
+Besides, the coefficient matrix in the virtual marker representation encodes spatial relationships among mesh vertices which allows the method to implicitly explore structure priors of human body.
+It achieves better mesh estimation results than the state-of-the-art methods and shows advanced generalization potential in spite of its simplicity.
+Acknowledgement
+This work was supported by MOST-2022ZD0114900 and NSFC-62061136001.
+Appendix
+We elaborate on the post-processing implementation of the virtual markers and provide additional experimental details and results.
+At last, we discuss data from human subjects and the potential societal impact.
+A.  Post-processing on Virtual Markers
 As described in Section 3.1, considering the left-right symmetric human body structure, we slightly adjust the learned virtual markers $\mathbf{Z}$ to be symmetric.
@@
 Figure 9: Illustration of the human mesh template $\mathbf{X}^{t}$ at the 3D coordinate system and a symmetric vertex pair $(\mathbf{X}^{t}_{i},\mathbf{X}^{t}_{j})$.
-Appendix B Experiments
+B.  Experiments
 In this section, we first add detailed descriptions for datasets and then provide more experimental results of our approach.
@@
 Following the previous works [32, 33, 24, 61], we use the train set of 3DPW to learn the model and evaluate on the test set.
-The same evaluation metrics as H3.6M are used.
-We further evaluate the performance of our probabilistic framework VMarker-Pro on two subsets of 3DPW, i.e.
-3DPW-OC and 3DPW-PC [14, 95] which have both object and person occlusion.
-In this setting, we do not use the 3DPW training set for a fair comparison following [96].
-MPI-INF-3DHP [97] is a 3D pose dataset with 3D GT pose annotations.
-Since this dataset does not provide 3D mesh annotations, following [6, 33], we only enforce supervision on the 3D skeletons (Eq.
-9) in mesh losses.
-UP-3D [15] is a wild 2D pose dataset with natural images.
+MPI-INF-3DHP [38]
+is a 3D pose dataset with 3D GT pose annotations. Since this dataset does not provide 3D mesh annotations, following [19, 25], we only enforce supervision on the 3D skeletons (Eq. (9)) in mesh losses.
+UP-3D [28]
+is a wild 2D pose dataset with natural images.
 The 3D poses and meshes are obtained by SMPLify [2].
@@
 Therefore we only use the 2D annotations to train the 3D virtual marker estimation network as in [47].
-COCO [98] is a large wild 2D pose dataset with natural images.
+COCO [34]
+is a large wild 2D pose dataset with natural images.
 Previous work [39] used SMPLify-X [41] to obtain pseudo SMPL mesh annotations but they are not accurate.
@@
 So we leverage the 2D annotations to train the virtual marker estimation network as in [47].
-SURREAL [27] is a large-scale synthetic dataset containing 6 million frames of synthetic humans.
+SURREAL [53]
+is a large-scale synthetic dataset containing 6 million frames of synthetic humans.
 The images are photo-realistic renderings of people under large variations in shape, texture, viewpoint, and body pose.
@@
 We use the same training split to train the model and evaluate the test split following [7].
-B.2 Implementation Details
+B.2  Implementation Details and Computation Resource
 Following common practice [19, 7, 39, 61, 26, 23, 33, 32], we conduct mix-training by using the above 2D and 3D datasets for experiments on the H3.6M and 3DPW datasets.
@@
 We implement the proposed method with PyTorch.
-The experiments for VMarker and VMarker-Pro are conducted on a Linux machine with 4 NVIDIA 16GB V100 GPUs and 2 NVIDIA A100-PCIE-80G GPUs, respectively.
-To train VMarker, the whole network is trained for 40 epochs with batch size 32 using Adam [100] optimizer.
-We set $\lambda_{3d},\lambda_{c},\lambda_{m},\lambda_{e}=100,0.1,0.1,0.1$ for each loss term.
-To train VMarker-Pro, the whole network is trained for 50 epochs with batch size 80 using Adam [100] optimizer.
-We set $\lambda_{diffusion}=1$ and keep the rest loss weights the same as training VMarker.
-For implementing LCN unit, we modify the original LCN network [77] by injecting a Cross-attention [82] module in each LCN layer.
-For more implementation details, please check https://github.com/ShirleyMaxx/VMarker-Pro.
+All the experiments are conducted on a Linux machine with 4 NVIDIA 16GB V100 GPUs.
+The whole network is trained for 40 epochs with batch size 32 using Adam [22] optimizer.
+We evaluate the model complexity in terms of FLOPs (G) and the number of model parameters in Table 7.
+Compared to the most recent state-of-the-art methods that directly regress all mesh vertices, such as I2L-MeshNet [39], METRO [32], and Mesh Graphormer [33], our approach with virtual marker representation reduces the computation overhead by a large margin while getting better estimation quality.
+The last column shows the MPVE errors on 3DPW test set for performance reference.
+[…]
+I2L-MeshNet [39] ECCV’20 28.7 141.2 110.1
+METRO [32] CVPR’21 153.0 397.5 88.2
+Mesh Graphormer [33] ICCV’21 48.8 180.6 87.7
+Ours 22.1 109.6 77.9
+Table 7: Computation overhead comparison with the recent state-of-the-art methods that directly regress all 3D vertices.
+The rightmost column shows the MPVE errors on the 3DPW test set for performance reference.
+Ours w/o $\mathcal{L}_{conf}$ w/o $\mathcal{L}_{pose}$ w/o $\mathcal{L}_{normal}$ w/o $\mathcal{L}_{edge}$
+[…]
+Table 8: MPVE errors on H3.6M [16] test set when ablating different loss terms.
+[…]
+None (Ours) 77.9 67.5 41.3
+2 Arms 79.2 $\uparrow$ 1.3 68.2 $\uparrow$ 0.7 42.2 $\uparrow$ 0.9
+2 Legs 78.3 $\uparrow$ 0.4 67.9 $\uparrow$ 0.4 41.7 $\uparrow$ 0.4
+Body 78.6 $\uparrow$ 0.7 68.0 $\uparrow$ 0.5 41.8 $\uparrow$ 0.5
+Random 78.7 $\uparrow$ 0.8 68.1 $\uparrow$ 0.6 41.9 $\uparrow$ 0.6
+Table 9: Results on 3DPW [54] test set when different parts of virtual markers (VM) are occluded.
 B.3  Additional Quantitative Results
+Figure 10: Meshes estimated by our approach on Internet images with challenging cases (complex poses or extreme body shapes).
 Different loss terms.
-Table XII reports the MPVE error on H3.6M [13] test set when ablating different loss terms.
+Table 8 reports the MPVE error on H3.6M [16] test set when ablating different loss terms.
 The confidence loss [17] is used to encourage the interpretability of the heatmaps to have a maxima response at the GT position.
@@
 If ablating the surface losses, MPVE increases a lot, which demonstrates the smoothing effect of these two terms.
+Robustness to occlusion.
+We report results when different virtual markers are occluded by a synthetic mask in Table 9.
+The errors are slightly larger than the original image (None), which validates the effectiveness of the locality of the virtual marker representation.
+Occluding arm regions results in a larger error increase.
+This may be because the arm has larger variations in the dataset.
 Comparison to fitting.
@@
 We then fit the mesh model parameters to these virtual markers.
-Table XIII shows the metrics of the fitted mesh on the SURREAL [27] test set.
+Table 10 shows the metrics of the fitted mesh on the SURREAL [53] test set.
 As we can see, the fitted mesh has a similar error as our regression ones which uses the interpolation matrix $\hat{\mathbf{A}}$, which validates the accuracy of the estimated virtual markers.
@@
 Figure 12 shows more qualitative comparisons with Pose2Mesh [7] on the SURREAL test set in which has diverse body shapes.
-The skeleton representation used in Pose2Mesh [1] loses the body shape information so the method can only recover mean shapes, regardless of whether the person is slim or stout.
+The skeleton representation used in Pose2Mesh loses the body shape information so the method [7] can only recover mean shapes.
+For example, in Figure 12 (d) (e), the estimated meshes of Pose2Mesh tend to have the average body shape and fail to estimate the real body shape, regardless of whether the person is slim or stout.
 This is caused by the limited skeleton representation bottleneck so that the model learns a mean shape for the whole training dataset implicitly.
-In contrast, both of our approaches VMarker and VMarker-Pro generate more accurate mesh estimation results with the help of virtual marker representation.
-TABLE XII: MPVE errors on H3.6M [13] test set when ablating different loss terms.
-VMarker w/o w/o w/o w/o
-$\mathcal{L}_{conf}$ $\mathcal{L}_{pose}$ $\mathcal{L}_{normal}$ $\mathcal{L}_{edge}$
-[…]
-TABLE XIII: Results on SURREAL [27] test set when the mesh is obtained by fitting to the estimated virtual markers.
+In contrast, our approach with virtual marker representation generates more accurate mesh estimation results.
 […]
      Fitting      44.6      34.8      29.5
-VMarker 44.7 36.9 28.9
-Figure 17: Meshes estimated by VMarker on Internet images with challenging cases (complex poses or extreme body shapes).
-Figure 18: Meshes estimated by VMarker-Pro on Internet images with challenging cases (complex poses or extreme body shapes).
-For each case in a row, we present three hypotheses, showing both a projected view and a side view in each column.
-The rightmost column shows the overlapped side views to unveil the subtle differences.
+     Ours      44.7      36.9      28.9
+Table 10: Results on SURREAL [53] test set when the mesh is obtained by fitting to the estimated virtual markers.
 Figure 13 shows more qualitative comparisons with Pose2Mesh [7] and METRO [32] on the 3DPW test set.
@@
 In addition, it is more robust to truncation and occlusion and aligns the image better.
-Figure 17 and 18 show more qualitative results of VMarker and VMarker-Pro on Internet images with challenging cases, respectively.
-These images include humans with diverse body shapes or complex poses.
-Our methods generalize well on the natural scenes.
-When occlusion happens, the probabilistic method VMarker-Pro could infer multiple reasonable poses, e.g.
-the different angles of the left arm of the running woman in the first case in Figure 18.
-Figure 19: Qualitative comparison between our methods and Pose2Mesh [1] on SURREAL test set [27].
-Our approaches generate more accurate mesh estimation results on images of diverse body shapes.
-We show 2 hypotheses generated by VMarker-Pro.
-Figure 20: Qualitative comparison between our method VMarker and Pose2Mesh [1], METRO [87] on 3DPW test set [14].
+Figure 11: Typical failure cases.
+(a) The right arm has inaccurate shape estimation due to the inaccurate virtual marker estimation around the arm when occluded.
+(b) Our method treats the lower arm of another person as its own due to occlusion.
+(c) Interpenetration around the right hand.
+Figure 14 shows more quality results of our approach on the 3DPW [54], H3.6M [16], MPI-INF-3DHP [38], and COCO [34] datasets.
+Figure 10 shows more qualitative results on Internet images with challenging cases, such as extreme body shapes or complex poses.
+Our method generalizes well on the natural scenes.
+Figure 11 shows typical failure cases, including inaccurate shape estimation and interpenetration, which are mainly caused by inaccurate 3D virtual marker estimation when occlusion occurs.
+But as expected, the rest body parts are barely affected due to the local and sparse property of the virtual marker.
+Figure 12: Qualitative comparison between our method and Pose2Mesh [7] on SURREAL test set [53]. Our approach generates more accurate mesh estimation results on images of diverse body shapes.
+Figure 13: Qualitative comparison between our method and Pose2Mesh [7], METRO [32] on 3DPW test set [54].
 Our approach is more robust to occlusion and truncation and generates more accurate mesh estimation results that align images well.
-Appendix C Human Subject Data
+Figure 14: Meshes estimated by our approach on images from the 3DPW [54] dataset (row 1-4), H3.6M [16] dataset (row 5), MPI-INF-3DHP [38] dataset (row 6), and COCO dataset (last 2 rows) [34].
+C.  Human Subject Data
 We use existing public datasets of human subjects in our experiments following their official licensing requirements.